A Bibliography of Publications about TEX

Nelson H. F. Beebe Center for Scientific Computing Department of Mathematics

University of Utah Salt Lake City, UT 84112

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)

22 July 1994 Version 1.153

Abstract

This bibliography records books and articles on TEX, , WEB, fonts, typography, indexing, and software related to these topics. ISBN and ISSN numbers are given, where available. A few troff-related items are included because of their strong overlap with TEX issues.

Title word cross-reference

3 [Mil86b].

1992/93 [Sch93].

22nd [Mil86a]. 23rd [Mil85]. 24th [Mil84, Mil86a]. 25th [Mil85]. 26th [Mil84].

5.25in [Bec91b].

7th [Zla92].

8th [Nad90].

ABC [DE89]. ABC-XYZapf [DE89]. Actes [And83a]. adaptation [BB89]. Adapting [RS85]. Adobe [Ado90]. Advanced [Her81, KD93]. Advances [Grä93c, Pie91]. affect [MBHL91]. Age [Whi88]. Agents [Sch88a]. akapicie [Ryć93]. Algorithm [LS85]. Algorithmes [H´84]. Alphabet [DE89, Knu79c]. Alphabetic [RS85]. American [Ano84]. amont [FPR84]. AMS [Ano90b, KZ89]. Analysis [DM91]. Analytical [And83b]. Anatomy [Law90]. Animation [Hay77]. Anleitung [WKS^{+}93a, WKS^{+}93b]. Annotated [SS91]. Anwender [Sch91b]. APL [Hoh88]. Applications [BU89, GR82, Vul92, Cla90]. approach [Won90, Gho83, GB83]. April [vV88, vV86]. Arabic [Mac]. Arbeitsheft [WKS^{+}93a]. archive [KW90]. Aren’t [Sch88a]. Art [Abi88, Bai70, Bai90, SP92, TW87]. Article [vdLL89]. Articles [Zap87]. Artifacts [DM91]. Arts [Arn72]. Aspects [vdVM88]. Assistants [Swa86]. Assisté [Ros84]. Associates [Man87]. Association [Ano84]. Astronomical [Hec92]. Astronomy [Hec92]. aufgabenorientierte [WKS^{+}93a, WKS^{+}93b]. Aufsätze [Tsc87]. August [Cla90]. Ausgewählte [Tsc87]. auteurs [Féd89]. Author [Bry88, Doo89, Sto92]. Authors [Ano87, Swa86, Kel85]. Automating [CH87]. available [KW90].

B [WKS^{+}93b]. Band [Kop94, Kopara, Koparb, Sch94]. Based [BU89, AP83]. Basic [Krö90]. Basics [vB93a, vB92b]. Beam [Cer85]. Beginner [Sam83, SL91, Sno92]. Beginners [KD93]. beim [Bin93]. benefits [Krö90]. better [PK82]. Bibliographical [Zap87]. Bibliography [SS91, And83b]. Bitmap [Cas93]. Bitmap-free [Cas93]. Bluebook [Ano91]. Book [Bai70, Bai90, Ano90a, PH92, Pop92, Ric78, SL91, Tsc91b, Wil83]. Bookmaking [Lee79]. Books [Ben91, SB90, Hur78]. Boston [Mil87]. Box [Bil89]. Braille [AJ88]. Breaking [KP81]. breaks [PK82]. Buch [Abd92, Abd93]. Buches [Tsc87]. Buchsatz [Sch89, Sch91b]. building [RKM88]. built [Gun85]. built-in [Gun85]. Buyers [Lab84].

C [GP89]. C-like [GP89]. Caching [FK82]. Cahiers [Cah ]. Calligraphy [Zap87]. Calling [Knu89b]. Cambridge [But81]. Capabilities [Mil90]. caractères [H´84]. Cartography [Her69]. Character [Nai91, Ada91, Tun80]. Characters [sT81, GH82, Li85, RS85]. Chel [MF86a]. Chemical [BU89]. CHI [Knu78b]. Chicago [Ano87, Ano82]. Chinese [GH82, HG83, Li85, sT81, Tun80].

Chiwriter [Cor86]. Choix [Féd89]. Choosing [PK82]. CMS [AMSV85]. Co [Bie90, Ryć93]. Code [Féd89]. Collection [BU89]. Colophon [Ado98]. Com [Lia83]. Com-pu-ter [Lia83]. Command [BB85]. Como [Luc85b]. comp.text.tex [com ]. Companion [GMS94]. Compatibles [otB86]. Complete [Zap87]. Completely [Ale92]. Composing [Ano88, LVC89]. Composition [Ale92, BE92, Rub88, Gho83]. Compositors [Har91b]. composizione [Bec91b]. Compress [GH82]. Computer [BU89, Hah91, Hah93, Her81, Her72, Knu86g, Rub88, WH76, Bia92, vdVM88, Knu80b]. Computerized [AJ88]. Computers [BU89, Bia92, BK85]. Computing [Nad90]. Con [Bec91b]. Concept [Knu81, Knu85, Knu82b]. concepts [FSS82a, FSS82b]. Concisely [Joh92]. Conference [BU89, Dés86, Luc85b, Mil84, Mil85, Mil86a, Mil87, Zla92, vV88, vV86, vdVM88, Cla90]. Congress [Pie91]. constructing [Wal87b]. Contemporary [Zap87]. contenente [Bec91b]. Contribution [WH76]. contributions [DE89]. Controller [Cer85]. Cookbook [Ado85b]. Coordinates [WH76]. Copy [But81, Swa86]. Copyrights [Sto92]. Cosmetic [OC90]. cost [Gun85, Cer85]. CPU [Doo90]. Craft [Wil83]. CWEB [KL93]. Czecholsovakia [Nad90]. czyli [Paw93].

dans [Bor90]. dansk [Lar89]. DANTE [Sch94, Sch93]. DANTE-TEX-Stammtisch [Sch93]. Darstellung [Kar92]. Decorative [MF87a]. defects [Ada91]. Delivery [CDMR85]. Descripción [PSH93]. Description [GB83]. Design [Rei88, BH86, Dai67, DE89, EPO , Eng87, HC88, Lee79, Nel78, Ric78, Rub88, Rug83, Whi88, Wil83, Zap87, Hur78, RKM88, Sou85a, Gho83, GB83, Knu79c, Tun80]. Design/Production/Editing [Lee79].

Designer [Har91a, RKM88]. Designing [Cra71, Sou85b]. Desktop [FN87, Hec92, Lun92, Bla90]. Detecting [DM91]. Developments [Zap87]. device [BB89, Utt85]. device-independent [BB89]. Devices [Hug88, Gun85]. Diagrams [Lev90]. Dictionary [Luc85a]. difference [Pap91]. Different [Ale92]. Diffusion [Knu87]. Digital [AH89, BD83, DM89, Kar87, Kar92, Knu87, Rub88, Uli87, GC88, Ika83]. Digitale [Kar92]. directional [BB89]. Directions [Knu79a]. disco [Bec91b]. Display [Ado88, Tuf83]. Dissemination [EPO ]. division [Dés84b]. dla [Paw93]. Document [Ala93, BBv88, Bro91, Dil93, FSS82a, FSS82b, Geh87, KD93, Lam85, vV88, vV86, Krö90, Won90, CDMR85]. Documentation [Dés86, KL93, Knu82e, Luc85b, Knu83c]. Documents [Hah91, Hah93, Ano88, Shu94, AP83, Eic90, And83a]. Does [AMS90]. dokumentów [Wyr92]. Donald [Gho83]. Dot [Knu87]. Drawing [BKW89, Gur94a, Lev90, LS85]. driver [Gun85, Bee87]. Druku [Tat88]. DTP [SP92]. Dublin [Mil84, Mil85, Mil86a]. Dutch [vdVM88]. Dvůr [Nad90]. DVI [Bee87, IM86b, IM86a]. Dzielenie [Koł87].

e.V [Sch94]. Ease [Abi88]. Editing [Bar86, Lee79, Ste81, Swa86, But81]. Editor [Bro91, Wal87b]. editoria [Bec91a]. Editorial [Swa86]. Editors [Bos86, Sch88a]. Education [BU89]. Eesti [Ibr92]. Effectiveness [DM91]. Egyptian [Hug88]. Einführung [Kat93, Kop88, Kop91b, Kop93, Kop94, Sch88b, Abd92, Abd93]. Einfürung [Koparb, Kop91a]. Einsatz [Bin93]. eLaTEX [Ala93]. Electronic [BH86, Ano87, EPO , GR82, Man87, Whi88, vV88]. Elementare [Tsc86]. Elements [Tob85]. elettronica [Bec91a]. Engineers [Bue90]. English [App85, JHR88]. Environment [Nad90, Hay87, Lub90, AMSV85].

Envisioning [Tuf90]. EPSILON [Knu78b]. Ergänzungen [Kopara]. Eric [Pop92]. Errors [DM91, Knu89c, Knu88]. Erweiterungen [Koparb]. Erweiterungsmöglichkeiten [Kop90, Kop91a]. Essays [Bar86]. Euler [KZ89, Sie85]. Euro [TEX a]. European [Dés86, Luc85b, Nad90, Zla92, Cla90]. EuroTEX [Zla92]. Evaluate [DM91]. Evaluating [Ada91]. Everyone [Hah91, Hah93, Pop92]. Example [Bor93, Bor92a]. Examples [GR82]. Exeter [Cla90]. experience [Krö90]. Experiments [AS85]. Explicative [Bar86]. Expos [Sch88a]. Expos/’e [Sch88a].

Factors [BM89]. Families [MF87b]. Family [BH86, Bee87, Knu80b, MF86a, MF87a, MF86b]. Faster [IM86a]. Features [Ano90b]. few [JHR88]. Feynman [Lev90]. Fifty [DE89]. File [Wal87a, Knu82c]. First [Luc85b, Mil84, Sam83]. Font [Ado90, Eng87, Knu81, Ada91, Dea92, GC88, FK82, HG83, Knu82b]. Fonts [vB93c, Grä93c, Kar92, Knu90a, WH76, vB92c, GC88, Gun85, Knu90b, MF86a, MF87a, MF86b, MF87b]. Fonttechnology [Kar94]. Form [Tsc91b, Ika83]. Formal [GB83]. Format [Ado90, Ric78]. Formate [Kar92]. Formats [Kar87, Kar92, Wal87a]. Formatter [Fuj94, KW89]. Formatting [Geh87, BB89, FSS82a, FSS82b, Noo83]. Formulae [Fuj94, Luc85a]. formulas [vdLCL89]. Fortgeschrittene [App88, App94]. FORTRAN [Her70, Her69]. Forward [Kar92]. Four [Mil90]. Fourth [Mil87]. Fragen [Tsc87]. FrameMaker [Dar93]. francais [Dés84b]. France [Dés86, vV88]. free [Cas93, Lub90]. French [Dés84a, Dés85]. Friendly [AMSV85]. Fun [Knu90a, Knu90b]. functional [Krö90].

Garmisch [BU89].

Garmisch-Partenkirchen [BU89]. Gebruikersgroep [NTG ]. Geleitwort [Kar92]. gemacht [WKS^{+}93a, WKS^{+}93b]. GENDRIV [Gun85]. Generalised [LS85]. Generalized [Bry88, Ano88]. Generation [Li85, H´84]. Gestalt [Tsc87]. Gestalten [Sie89]. Gestaltungsbuch [Rud88]. Getting [IM86a]. Gourmet [Spi86a, Spi90]. GRAB [Les88]. Grade [Sam83]. Grand [Knu90a, Knu89b, Knu90b]. Graphic [Arn72, Whi88, WH76]. Graphical [Mol93]. Graphics [Lev90, Rah89, Whi83, Wal87b, Van80]. Grayscale [Nai91]. Great [Whi90]. grid [Hur78]. Group [Cor86]. Guida [Bec91a, Bia92]. Guide [Ala93, Bia92, BR93, Bry88, Ano87, Hah91, Hah93, Hol89, KD93, Lam85, Lee79, Spi86a, Spi90, Bla90, Rol93]. Guidelines [SvdL89]. GUTenberg [Cah ].

Hahn [Pop92]. Halftones [Knu87]. Halftoning [Uli87]. Handbook [Arn72, BR93, Har91a, Hig93, But81]. Happen [AMS90]. hax [TEX b]. held [Mil84, Mil85, Mil86a]. Hermann [DE89, Kar92]. Hershey’s [WH76]. Herwijnen [Pop92]. Hieroglyphs [Hug88]. hints [Dea92]. HITAC [Hay87]. Horstman [Cor86]. Hortatory [Bar86]. Houses [Sch88a]. Hrsg [Sch94]. Hudson [McL80]. Human [BM89, vdVM88]. Human-computer [vdVM88]. Hy [Lia83]. Hy-phen-a-tion [Lia83]. Hyphenation [App85, Dés85].

I. [Man87]. IBM [otB86]. ICL [Lei85]. II [Mil85]. III [Mil86a]. IKARUS [Ika83]. Illuminating [TW87]. Illustrated [Lee79]. Illustrations [Zap87]. image [Ada91]. Images [Ros84]. Imaging [AH89]. impact [Pie91]. Impatient [wKBH90]. impaziente [wKBH91]. Implementations [Grä93c]. Implementing [Lei85]. Importance [DM89, Krö90]. Improved [Nai91]. inaczej [Wiȩ92]. Inc. [Mil86b]. independent [BB89, Utt85]. Index [CH88, CH87, Che90]. Indexes [BK88, Les88]. Indian [Gho83]. Industrial [Wil83]. Information [Tuf90, Tuf83]. Informatique [And85]. initial [Krö90]. Ink [Arn72]. Inleiding [Sch90]. Inputting [sT81, FPR85]. InScript [GP89]. Insider [Sch88a]. Intégrant [Ros84]. Integrated [BH86]. Interaction [vdVM88]. Interactive [AMSV85, AGQ85]. interchange [Krö90]. Interface [ASW89, Nad90]. interfaces [LVC89]. International [BU89, Mil84, Mil85, Mil86a, Mil87, Pie91, Sch91a, vV88, vV86, Par90]. InterViews [LVC89]. Introduction [Grä93a, KS90, Rub88, GC88, Urb86]. Inverted [Les88]. IS&T [Pie91]. ISO [Krö90, vH85]. ISO/STPL [vH85]. issues [FSS82a, FSS82b]. Italian [AP83]. Italy [Luc85b]. IV [Her69, Mil87].

Jahrbuch [Sch93]. Jane [Pop92]. janvier [And85]. Japanese [Ano90a]. jest [Bie90]. Journal [SvdL89, Win88, POS]. Journées [And83a]. Joy [Spi86a, Spi90]. June [Dés86].

Keyboard [FPR85]. Know [Sch88a]. Knuth [Gho83]. komendach [Dyl94]. Komödie [DAN ]. Komputerowy [Dyl94].

Ładnego [Tat88]. LAMS-TEX [Spi89]. Language [Rei88, Ado85a, Ado85b, Bry88, Ano88, Gou86, Bie88, Tun80, Van80, POS]. Languages [RS85]. Laser [Cer85]. LaserJet [Pap90]. LaTEX-Buch [Abd92, Abd93]. Latin [RS85]. laurea [Bec91b]. layout [Eic90]. LCCD [Tun80]. Learned [Knu83a]. lecons [Spi86b]. Lectures [BU89, Zap87]. Lehrbuch [Sch91b]. leicht [WKS^{+}93a, WKS^{+}93b]. Lernen [WKS^{+}93a, WKS^{+}93b]. Lessons [Knu83a]. Letter [Knu80a, Knu80c, vdLL89]. Letterform [Rug83]. Lettershape [GB83]. Library [IM86b, IM86a]. Lifetree [Mil86b]. like [Fat87, GP89]. Lilac [Bro91]. Line [BR93, Dil93, PK82]. Lines [KP81]. List [Zap87]. Literacy [BH86]. Literate [Ben86, Gur94a, Knu84b, Knu92, Knu83b, Sew89, SS91]. Livre [Ser89, Féd89]. Logical [Eic90]. logiciel [Dés84b, FPR84, Koł88b]. Low [Les88, Gun85, Cer85]. Lucida [BH86]. ludzi [Paw93].

Machine [Nad90]. Macintosh [Mol93]. macro [Spi86a, Spi90]. macroistruzioni [Bec91b]. Macros [vB93d, vB92d]. MACSYMA [Fat87]. MACSYMA-like [Fat87]. magazines [Hur78]. Mai [And83a]. make [Par90]. MakeIndex [Sch91a]. Makeup [KW89]. Making [Wal94]. Man [Nad90]. Man-Machine [Nad90]. Manifesto [Sie86]. Manipulation [vV88, vV86, And83b, And83a]. Manual [Ano84, Ado85a, Ano82, HC88, Lab84, Lam85, McL80, Sam83, Spi91, Spi86b, Spi85, WKS^{+}93b]. Manuscript [AMS90]. Manuscripts [Ano87]. Mark [AJ88]. Mark-up [AJ88]. Markup [Bry88, Ano88]. master [KW90]. Mastering [Whi83]. Matching [GH82]. Matematisk [Mor91]. Math [vB93c, Grä93a, vB92c, AGQ85]. Mathematical [Bor93, Hig93, Knu79b, KLR89, Knu78a, Mor91, Pal86, vdLCL89]. Mathematics [KZ89, Swa86]. mathématiques [FPR84, Roy84]. May [Luc85b]. Media [GR82]. Meta [Knu81, HG83, Knu82b]. Meta-Font [Knu81, HG83, Knu82b]. Metafont [Jac93, Knu79c, Sch94, Tob85, Li85, Sou85b, KW90]. book [Knu86e]. Metafonte [Knu85]. Metamarks [Bil89]. method [Ada91]. Methods [Cla90, Wil83]. MF [MF86a, MF87a, MF86b, MF87b]. Mitteilungen [Tsc86]. Mode [AGQ85]. Models [Ric78]. Modern [Hug88, Knu86g, Vul92, Knu80b]. modular [Hur78]. mots [Dés84b]. MS [Bec91b].

MS-DOS [Bec91b]. music [Gou86].

Mutation [DM91].

naturvidenskabelig [Mor91].

Nederlandstalige [NTG ]. Neue [Tsc91a]. News [Ado98]. Newspaper [Har91a]. newspapers [Hur78]. Newswriting [GR82]. Nice [vV88]. Nicely [BKW89]. nieco [Ryć93]. niepraktyczne [Jac93]. Non [Pie91, JHR88, App85, LS85]. non-English [JHR88, App85]. Non-impact [Pie91]. Non-Parametric [LS85]. note [Bec91b]. Notes [Shu94, Zap87]. Nottingham [vV86].

o [Ryć93]. Observatory [Hec92]. Occidental [WH76]. October [Hec92, Mil84, Mil85, Mil86a, Mil87, Pie91]. Offizin [Sch94]. On-Line [BR93]. Optimal [FK82]. ordinateur [Dés84b]. Original [Eng87]. Origination [EPO ]. Outline [Dea92]. Output [vB93b, Fat87, KW89, vB92e, Pap90, IM86a, Utt85]. Overhead [Les88]. Oxford [Har91b].

P [Man87]. Package [Cas91, Spi86a, Spi90]. Page [Ale92, KW89]. Pages [Whi90]. Pandora [Bil89]. Paper [AMS90, Arn72]. Papers [BU89]. Paragraphs [vB93c, vB92c, KP81]. Parametric [LS85]. parochial [JHR88]. Part [AMS90]. Partenkirchen [BU89]. pasos [Gal92]. PC [otB86, Wal87a]. PCTEX [Spi86b, Spi85]. perfection [Bor90]. Permuted [Che90]. Perq [Lei85]. personal [Bia92, DE89]. petit [Ser89]. Phase [DM89]. phen [Lia83]. Philological [Won92]. Philosophy [Zap87]. Physics [Nad90]. Plain [Cla92]. podrecznik [Wyr92]. Poles [JHR88]. Polish [Bie88]. polonais [Koł88b]. polskich [Koł87]. polskie [Wiȩ92]. Popular [Wal87a]. Porting [Lub90]. Portland [Pie91]. Postprocessing [KW89]. PostScript [Grä93c, Hol89, Hol87, GP89, Pap90]. Practical [Lab84, Pop92, Shu94, vH90].

Practice [vB93a, vB93c, vB93d, vB93b, Wil83, vB92a, vB92b, vB92e, vB92c, vB92d, Mor91, RKM88]. praksis [Mor91]. praktyczne [Jac93]. pratique [Rol93]. praxisorientierte [Abd92, Abd93]. Pre [Sie86]. predicting [Ada91]. preface [Gho83]. Preliminary [Bil89]. Premiers [Urb91]. Preparation [Ala93, CH88, CH87, Her70, KD93, Lam85]. Preparing [Ano87, Shu94]. preprocessor [GP89]. Presentation [Nai91]. Presented [BU89]. Press [Har91b]. Primer [Cla92, SK93, Dea92]. primeros [Gal92]. Principles [GR82]. printer [Pap90, Cer85]. Printing [Ben91, BK88, Hug88, Pie91, Gou86]. Problem [sT81, Sou85a]. Proceedings [Dés86, Luc85b, Mil84, Mil85, Mil86a, Mil87, Zla92, vV88, vV86, Cla90, Nad90]. Processing [Abi88, CH88, Dil93, Mil84, Mil85, Mil86a, Mil87, Mil86b, Pal86, vV86, BBv88, Won90]. Processor [Lab84]. Processors [Mil90, Mil92]. Produce [Win88]. Production [DM91, Lee79, AP83, Hur78, Ros84]. Profesjonalny [Bor92b]. Professional [DE89]. Professioneller [Sch89, Sch91b]. professionels [Féd89]. Program [Rei88, Knu86d, Knu86f, Sew89]. Programmer [Hol89, Wal87a]. Programming [Ben86, Gur94a, Her69, Hol87, Knu84b, Knu92, Knu83b, Sew89, SS91]. Programs [BM89, Mil86b, Lub90]. Project [Sie85]. Proofreading [Swa86]. PROTEXT [Mil84, Mil85, Mil86a, Mil87]. Przewodnik [Dyl94, Wyr92]. przygotowywania [Wyr92]. przykładami [Dyl94]. Psychological [Ano84]. Psychonomic [vdVM88]. pu [Lia83]. Publication [Ano84, Nel78, Ado98]. Publiceren [dBvLV88]. publikacji [Bor92b]. Published [AMS90]. Publisher [Gil89]. Publishers [Ano87]. Publishing [Bai70, Bai90, Bar86, EPO , FN87, Hec92, HC88, Lun92, Man87, Sch88a, vV88, Bla90].

quality [Ada91]. Quantitative [Tuf83].

Rôle [Utt85]. rapporter [Mor91]. Raster [AH89]. rasterization [Dea92]. rather [SP92]. Rationalized [FPR85]. readability [MBHL91]. Readable [BM89]. Readers [Har91b]. Real [Rot88]. REDUCE [ASW89]. Reference [Ado88, Ado85a, Eij92, Hah91, Hah93, Hol89, Lam85, Wal87a]. règles [Féd89]. remarks [JHR88]. Rennes [And83a]. Repertory [WH76]. Report [Sch88a]. Reports [Her70, Mor91]. Representation [Kar92]. requirements [RKM88]. Research [BU89, Mil86b]. resources [GC88]. Responsibilities [Sto92]. Review [Mil86b]. reviews [PH92, Pop92]. Revision [Bee87]. Riad [BK85]. Risø [Mor91]. Risø-rapporter [Mor91]. Roman [MF86b]. Routine [Lev90]. Routines [vB93b, vB92e]. Rules [Har91b]. run [Dés84a].

S [Knu80a, Knu80c]. S.A.PR.I.S.T.I

[Ros84]. Satzkunst [SP92]. School [Nad90]. Schriften [Kar92]. Schriftenreihe [Sch94]. Science [Bai70, Bai90]. Sciences [Hec92, Hig93]. Scientific [Dés86, Her72, Luc85b, Mil86b, Mil92, Nad90, Mor91, CDMR85, Mil86b]. scientifica [Bec91b]. Scientists [Bue90, SK93]. script [Mac]. Scripts [Gho83]. Scriptwriting [Hay77]. SDDS [CDMR85]. Second [Dés86, Mil85]. Selected [Zap87, DE89]. September [Nad90]. Seventh [Pie91]. SGML [Bar89a, Bry88, vH90, vdLCL89, Pop92]. Shapes [Bil89]. Sharp [Man87]. Simplement [Ros84]. Simplified [Grä93a]. Single [Doo90]. sistema [Bec91a]. sizes [MBHL91]. Skalský [Nad90]. Skills [FN87]. skład [Bor92b, Dyl94]. Slavic [MF87b]. small [MBHL91]. Society [vdVM88]. Software [Cor86, DM91, Mil86b, Win88, Wal87a]. Some [Li85]. Sonderheft [Tsc86]. sophisticated [Lub90]. sources [KW90]. Space [Hec92]. spec [Whi87]. Spectra [DM89]. Splines [LS85]. Springer [Bin93]. Springer-Verlag [Bin93]. środowiskach [Dyl94]. Stammtisch [Sch93]. Standard [Bry88, Krö90, vH85]. Standards [AJ88, HC88, Krö90]. Stanford [Sie85]. Stanford’s [KW90]. Starting [Doo93]. statt [SP92]. STET [Bos86]. stile [Bec91b]. Storage [Les88]. STPL [vH85]. Strasbourg [Dés86, Hec92]. STRATEC [FPR84, FPR85]. String [GH82]. Structural [Fuj94]. Structured [KL93, Knu82e, Noo83, Won90, Knu83c]. structures [Eic90]. studies [Bil89]. Style [Ano82, HC88, OC90, SvdL89, Tob85]. Summary [BB85]. Summer [Nad90]. Supposed [Sch88a]. Survey [Rah89, FSS82a, FSS82b]. Symbols [WH76]. Synthesis [Spi89]. System [Ado88, Geh87, Her72, Her70, KL93, Knu82e, Lam85, Rub88, Wyr92, AP83, Hur78, CDMR85, Knu78b, Knu79c, Knu83c]. Système [Ros84]. systemie [Dyl94, Koł87]. Systems [Mil84, Mil85, Mil86a, Mil87, Rug83, Fat87, FSS82a, FSS82b, Ado98]. Sztuka [Lao86].

T [Mil86b]. Tables [vB93b, WH76]. Tangled [Mor93]. TAU [Knu78b]. TCI [Mil86b]. Teach [Chu92]. teaching [AS85]. Technical [otB86, Shu94, Ste81, Knu78b]. Techniques [Dil93, Nad90, WH76]. Technologies [Pie91]. technology [Dea92]. Teil [WKS^{+}93a, WKS^{+}93b]. tekstu [Dyl94]. Templates [vdLL89]. ter [Lia83]. tesi [Bec91b]. test [Knu84c, Knu82c, Knu86a]. TEX [KW90]. TEX-Euro [TEX a]. TEX-REDUCE [ASW89]. TEX/LaTEX [Fuj94]. TEX/ [KW90]. TEX/WEB [Roy84]. TEX82 [Knu82a]. TEX88 [Cla90]. TEXbook [Knu84a, Knu89a, Knu86c]. TEXHelp [BR93]. TEXiraamat [Ibr92]. TEXline [TEX c]. TEXmag [TEX d]. TEXnician [Eij92]. TEXnicolor [Ano90a]. TEXnika [Tat88]. TEXnische [DAN ]. TEXowania [Lao86]. Text [Abi88, Fuj94, KW89, Mil84, Mil85, Mil86a, Mil87, Pal86, Ric78, Ste81, vV86, And83b, BB89, Noo83, Gho83, Knu78b]. texte [Bor90]. Textes [Ros84, FPR84, Koł88b, Roy84]. TEXtures [Ale92]. TEXware [Knu86b, Knu82d]. Thames [McL80]. Third [Mil86a, Cla90]. Three [Mil92]. tion [Lia83]. Tips [Dil93, Shu94]. Tokens [vB93d, vB92d]. Tools [BK88, RKM88]. Topic [Eij92]. Topography [AH89]. torture [Knu84c, Knu86a]. Tour [Mol93]. Trade [Bos86]. traitement [Bor90, FPR84, Koł88b, Roy84]. Transporting [Hay87]. Trees [BKW89]. tri [BB89]. tri-directional [BB89]. Tricks [Bos86]. triroff [BB89]. troff [BB89]. Tutorial [Ado85b, Hah91, Hah93]. Two [Bro91, Mil86b]. Two-View [Bro91]. Type [Ado90, Cra71, Dai67, DM89, Lab84, Lun92, Rub88, SG74, Swa86, WH76, Zap87, Sou85a, Whi87, Gho83, GB83]. Typeface [KZ89, Law90, MBHL91]. Typefaces [Bla90, Kar87, Knu86g, Sou85b, SB90, Zap87, Ika83, Knu80b]. Types [BH86]. Typeset [Lab84]. Typesetters [Lab84]. Typesetting [AJ88, BE92, Cas91, Doo89, Fuj94, Geh87, Gil89, Hah91, Hah93, Hoh88, Kel85, Knu79a, Mol93, Spi86a, Spi90, WH76, Won92, Pap91, Sie86, Van80]. typografi [Mor91]. Typographic [Her70, OC90, GC88, MBHL91]. Typographie [And85, Rud88, Tsc91a, Tsc86]. Typographique [Féd89]. Typographische [Tsc86]. Typographisches [Sie89]. Typography [BM89, BD83, Her81, Her72, Her69, Knu79b, Knu78a, McL80, Rub88, SP92, vV88, Mor91, Dés85].

użytkownika [Wyr92]. UKTEX [UKT ]. Understanding [Hol87, Lun92]. University [Har91b, vV86]. UNIX [Geh87, Lub90]. USA [Pie91, Mil87]. usage [Féd89]. Use [Bar89a, DM91, Nai91, Bia92, RS85, Dés85, Sie86]. used [Win88]. User [BR93, Lam85, Bla90, LVC89, RKM88, AMSV85]. User-Friendly [AMSV85]. Users [KD93]. Uses [Cla90]. Using [Chu92, Eng87, Grä93a, GH82, Hah91, Hah93, Man87, Tri88, Gun85, JHR88, IM86b, IM86a]. uso [Bia92].

Väike [Ibr92]. variation [MBHL91]. Vein [Ale92]. ver.0.93 [FPR84]. ver.2.3 [Knu83c]. Verlag [Bin93]. Version [KL93, Knu82e, Sch91a, Knu82d, Mil86b]. verze [LU92]. View [AJ88, Bro91, Doo89, Gil89]. Vieweg [Abd92, Abd93]. VIIth [BU89]. Virtual [Knu90a, Knu90b]. Visual [Tuf83]. VM [AMSV85]. VM/CMS [AMSV85]. Volkswriter [Mil86b]. VOS3 [Hay87]. VTEX [Cas91].

w [Dyl94, Koł87]. Wat [EP90]. Weaving [Sew89]. WEB [Knu83c, Knu82e, Roy84]. were [BU89]. Wider [Sie86]. Wizard [Spi91]. Wizards [Knu89b, Knu90a, Knu90b]. Word [Lab84, Lia83, Mil86b, Mil90, Mil92]. Word-Processing [Mil86b]. Wordprocessors [otB86]. Words [SG74, App85]. Work [Wal94]. Workshops [Chu92]. World [Rot88]. wraz [Dyl94]. Writers [Bos86]. Writing [Bar86, Gur94b, Hig93, KLR89, Sch88a]. Wuppertal [Sch93]. wyrazów [Koł87].

Xerox [HC88]. XYZapf [DE89].

Years [DE89].

z [Dyl94]. Zapf [DE89, Kar92]. zastosowania [Jac93]. znaki [Wiȩ92].

References

Abdelhamid:VLB92

[Abd92] Rames Abdelhamid. Das Vieweg LaTEX-Buch: Eine praxisorientierte Einführung. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1992. ISBN 3-528-05145-0. 200 pp.

Abdelhamid:VLB93

[Abd93] Rames Abdelhamid. Das Vieweg LaTEX-Buch: Eine praxisorientierte Einführung. Friedrich Vieweg und Sohn, Braunschweig, Germany, second edition, 1993. ISBN 3-528-15145-5. xvi + 169 pp.

Abikoff:MI-8-3-64

[Abi86] William Abikoff. TEX. The Mathematical Intelligencer, 8(3): 64–76, 1986.

Abikoff:TEA88

[Abi88] William Abikoff. TEX: The ease and art of text processing. Abacus, 5(4):10–29, summer 1988.

Adams:NIPT-353

[Ada91] Debra A. Adams. Evaluating font quality: A method of predicting character image defects. In Pietrowski [Pie91], pages 353–

Adobe:colophon

[Ado98] Adobe Systems Incorporated.

Colophon—Adobe Systems News Publication, 198?

Adobe:PLR85

[Ado85a] Adobe Systems Incorporated.

PostScript Language Reference Manual. Addison-Wesley, Reading, MA, USA, 1985. ISBN 0-201-10174-2. ix + 321 pp. LCCN QA76.73.P67 A33 1985.

Adobe:PLT85

[Ado85b] Adobe Systems Incorporated.

PostScript Language Tutorial and Cookbook. Addison-Wes-ley, Reading, MA, USA, 1985. ISBN 0-201-10179-3. x + 243 pp. LCCN QA76.73.P67 A34 1985.

Adobe:DPS88

[Ado88] Adobe Systems, Inc. The Display PostScript System Reference, October 10 1988.

Adobe:AT190

[Ado90] Adobe Systems, Inc. Adobe Type 1 Font Format, March 1990.

Andre:TEX85-79

[AGQ85] Jacques André, Yann Grundt, and Vincent Quint. Towards an Interactive Math Mode in TEX. In Lucarella [Luc85b], pages 79– 92. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Andre:RID89

[AH89] Jacques André and Roger Hersch, editors. Raster Imaging and Digital Topography. Cambridge University Press, New York, 1989. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Arrabito:EPODD-1-2-117

[AJ88] R. Arrabito and H. Jürgensen. Computerized Braille typesetting: Another view of markup standards. Electronic Publishing—Origination, Dissemination, and Design, 1(2):117–132, September 1988.

Alamnehe:EDP93

[Ala93] Abass Alamnehe. eLaTEX: Document Preparation Guide. EthiO Systems Co., P.O. Box 36921, Houston, TX 77236, USA, 1993. xvi + 230 pp. This book is typeset in eLaTEX, an Ethiopicbased version of LaTEX, and is entirely in the Amharic language, including many of the control sequences.

Alexander:SEYBOLD-7-1-3

[Ale92] George Alexander. TEX and TEXtures: Page composition in a completely different vein. The Seybold Report on Desktop Publishing, 7(1):3–7, September 4 1992.

AMSCSD:NAMS-37-8-1027

[AMS90] AMS Composition Services Department. From manuscript to published paper: How does it happen? part 2. Notices of the American Mathematical Society, 37(8):1027–1028, October 1990.

Agostini:TEX85-117

[AMSV85] M. Agostini, V. Matano, M. Schaerf, and M. Vascotto. An Interactive User-Friendly TEX in VM/CMS Environment. In Lucarella [Luc85b], pages 117–132. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Andre:AJM83

[And83a] Jacques André, editor. Actes des Journées sur la Manipulation de Documents, Rennes 4–6 Mai 1983. INRIA/IRISA, Rennes, France, 1983.

Andre:TSI-1-5

[And83b] Jacques André. Analytical bibliography on text manipulation. Technology and Science of Informatics, 1(5), 1983.

Andre:INRIA85

[And85] Jacques André, editor. Typographie et Informatique, 21– 25 janvier 1985. INRIA/IRISA, Rennes, France, 1985.

Chicago:CMS82

[Ano82] Anonymous. The Chicago Manual of Style. University of Chicago Press, Chicago, IL, USA, 13th edition, 1982. ISBN 0-226-10390-0. ix + 737 pp. LCCN Z253 .U69 1982. UK£30.

APA:PMA84

[Ano84] Anonymous. Publication Manual of the American Psychological Association. American Psychological Association, Washington, DC, USA, third edition, 1984. ISBN 0-912704-57-8. 208 pp. LCCN BF76.7 .P83 1983.

CGP87

[Ano87] Anonymous. Chicago Guide to Preparing Electronic Manuscripts (For Authors and Publishers). University of Chicago Press, Chicago, IL, USA, 1987. ISBN 0-226-10392-7 (hardcover), 0-226-10393-5 (paperback). xi + 143 pp. LCCN Z286.E43 U54 1987. Short bibliography, examples, appendices, index.

IBM:CDG88

[Ano88] Anonymous. Composing Documents with the Generalized Markup Language. Number S544-3421. IBM Corporation, San Jose, CA, USA, second edition, March 1988.

Japanese-texnicolor

[Ano90a] Anonymous. The Japanese TEXnicolor Book. ASCII Corporation, Tokyo, Japan, 1990. ISBN 4-7561-0405-3. Japanese; contents include chapters on the history of TEX and of the Japanese TEX extension, file formats (fonts and .dvi), and device drivers.

AMS:NAMS-37-2-143

[Ano90b] Anonymous. New TEX features from AMS. Notices of the American Mathematical Society, 37(2): 143–144, February 1990.

HLRA:91

[Ano91] Anonymous. The Bluebook. Harvard Law Review Association, Cambridge, MA, USA, 15th edition, 1991. xvii + 343 pp. LCCN KF245 .U5 1991.

Aiello:RIV-INF-13-3-261

[AP83] L. Aiello and S. Pavan. A system for the production of documents based on TEX (in italian). Riv. Inf. (Milan), 13(3):261–270, July 1983.

Appelt:TEX85-61

[App85] Wolfgang Appelt. The Hyphenation of Non-English Words with TEX. In Lucarella [Luc85b], pages 61–65. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Appelt:TF88

[App88] Wolfgang Appelt. TEX für Fortgeschrittene. Addison-Wesley, Reading, MA, USA, 1988. ISBN 3-89319-115-1. x + 179 pp.

Appelt:TF94

[App94] Wolfgang Appelt. TEX für Fortgeschrittene. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 3-89319-695-1. xii + 183 pp.

Arnold:IP72

[Arn72] Edmund C. Arnold. Ink on Paper 2: a Handbook of the Graphic Arts. Harper & Row, Publishers, New York, NY, USA, 1972. ISBN 0-06-010131-8. 374 pp. LCCN Z244 .A83 1972. Chapter bibliographies, sidebars, glossary, index, pp. 375.

Andre:TEX85-141

[AS85] Jacques André and Richard Southall. Experiments in teaching . In Lucarella [Luc85b], pages 141–153. ISBN 0-201-13399-7. LCCN Q224.E97 1985. Antweiler:SIGSAM-23-2-26

[ASW89] Werner Antweiler, Andreas Strotmann, and Volker Winkelmann. A TEX-REDUCE interface. SIGSAM Bulletin, 23(2): 26–33, April 1989.

Bailey:ASB70

[Bai70] Herbert S. Bailey, Jr. The Art and Science of Book Publishing. Harper & Row, Publishers, New York, NY, USA, 1970. xii + 216 pp. LCCN Z278 .B15. Bibliography, tables, figures, appendix, index, pp. 217.

Bailey:ASB90

[Bai90] Herbert S. Bailey, Jr. The Art and Science of Book Publishing. Ohio University Press, Athens, OH, USA, third edition, 1990. ISBN 0-8214-0970-0. xviii + 216 pp. LCCN Z278 .B15 1990.

Barzun:WEP86

[Bar86]  Jacques Barzun.  On Writ-

ing, Editing, and Publishing: Essays, Explicative and Hortatory. University of Chicago Press, Chicago, IL, USA, 1986. ISBN 0-226-03857-2 (hardcover), 0-226-03858-0 (paperback). xi + 148 pp. LCCN PN149 .B295 1986.

Barron:EPODD-2-1-3

[Bar89a] D. Barron. Why use SGML? Electronic Publishing—Origination, Dissemination, and Design, 2(1):3–24, April 1989.

Barwise:NAMS-36-3-241

[Bar89b] Jon Barwise. TEX, LaTEX, and AMS-TEX. Notices of the American Mathematical Society, 36(3): 241–243, March 1989.

Botway:LCS-85

[BB85] Lloyd Botway and Chris Biemesderfer. LaTEX Command Summary (TEX 1.3, LaTEX 2.08). Prepared for the Space Telescope Institute, Baltimore, MD; reprinted with permission and distributed by TUG, 1985.

Becker:EPODD-2-3-119

[BB89] Zeev Becker and Daniel Berry.

triroff, an adaptation of the device-independent troff for formatting tri-directional text. Electronic Publishing—Origination, Dissemination, and Design, 2(3):119–142, October 1989.

Bouma:DP88

[BBv88] Louis G. Bouma, Jeroen Bruijning, and Johannes C. van Vliet. Document processing. In van der Veer and Mulder [vdVM88], pages 144–169. ISBN 0-387-18901-7.

Bigelow:SA-249-2-106

[BD83] Charles Bigelow and Donald Day.

Digital typography. Scientific American, 249(2):106–119, August 1983.

Barlow:TC92

[BE92] Geoff Barlow and Simon Eccles. Typesetting and Composition. Blueprint, London, UK, second edition, 1992. ISBN 0-948905-72-7. 274 pp.

Beccari:LGS91

[Bec91a] Claudio Beccari. LaTEX—Guida a un sistema di editoria elettronica. Editore Ulrico Hoepli, Ulrico Hoepli Editore S.p.A., via Hoepli 5, 20121 Milano, Italy, 1991. ISBN 88-203-1931-4. ix + 399 pp. [From the author]: The book is the first and only one in Italian at this moment [18-Nov-1991]. Although it contains an introductory chapter for the beginner, its main purpose is to give the reader advanced information on LaTEX (and partially on TEX) in order to prepare sophisticated macros, style files, and to make a clever use of fonts. Here is a reduced table of contents (translated to English): Chapter 1 - Introduction

2 - Introductory LaTEX 3 - General commands

for text mode

4 - Mathematical commands

5 - Environments 6 - Figures and tables 7 - Macros 8 - Typical documents

and style files

9 - Fonts Appendix A - TEX Users Group

B - Keyboards C - Font tables D - Hyphenation pat-

terns for Italian Bibliography Index The publisher received cameraready pages that were prepared by the author with a 300-dpi laser printer. The page style dimensions were suitably enlarged while the photo plates were reduced so as to minimize the granulosity of the signs. In practice the book was prepared with 12pt option and reduced by the photografic ratio of 0.8 so that the final real font size is 9.6pt. The whole book was prepared with the standard Computer Roman fonts on purpose in order to show the potential user what he can get witha standard TEX/LaTEX kit. The simple pictures were prepared with the “picture” environment. The cover and the title page were designed by a graphic expert of the publishing company.

Beccari:TLS91

[Bec91b] Claudio Beccari. La tesi di laurea scientifica—Con note per la composizione in LaTEX—disco MS-DOS da 5.25in contenente macroistruzioni di stile. Editore Ulrico Hoepli, Ulrico Hoepli Editore S.p.A., via Hoepli 5, 20121 Milano, Italy, 1991. ISBN 88-203-1932-2. vii + 116 pp. The book was typeset by the author and printed with a laser printer having a resolution of 300 dpi. The pages were then photographically reduced by a factor of 0.8 to the final size.. The fonts used are those of the standard TEX/LaTEX set, that is Computer Modern, with a default normal size of 12pt that the photographic reduction made equivalent to approximately 10pt. The book is a set of instructions for the preparation of the thesis (Master and PhD levels) according to the Italian State Universities regulations; it deals mostly with literary style but it also includes extensive tables of scientific symbols and nomenclature and many references to Standards regulations. The final chapter deal with the typesetting of the thesis with a LaTEX style, TOPTESI, whose source is recorded in a floppy disk that comes with the book.

Beebe:dvidriver207

[Bee87] Nelson H. F. Beebe. A TEX DVI Driver Family — (Revision 2.07). University of Utah, Salt Lake City, UT 84112, USA, April 1987.

Bentley:CACM-29-5-364

[Ben86] Jon Bentley. Literate programming. Communications of the Association for Computing Machinery, 29(5):364–369, 1986.

Bennett:BP91

[Ben91] Paul A Bennett, editor. Books and Printing. Biel, Savannah, Ga, 1991. ISBN 0-913720-72-0. 417 pp. LCCN Z116.A3 B58 1991. first published in 1951.

Bigelow:TPDM86-1

[BH86] Charles Bigelow and Kris Holmes. The design of Lucida: an integrated family of types for electronic literacy. In van Vliet [vV86], pages 1–17. ISBN 0-521-32592-7. LCCN QA76.9.T48 T481 1986.

Bianco:GUP92

[Bia92] Fiorella Bianco. TEX—Guida all’uso su personal computer (Guide for the use on personal computers). Pitagora Editrice, Via del legatore, 3, Bologna, Italy, 1992. ISBN 88-371-0583-5. ix + 157 pp. Prepared with Plain TEX and book macros written by the author. Final copy produced on a 300-dpi laser printer at magnification 1.2, then reduced for printing.

Bien:TEXLINE-89

[Bie88] Janusz S. Bień. Polish Language and TEX. TEXline, (8):2, 1988.

Bien:WM-29-1-132

[Bie90] Janusz S. Bień. Co to jest TEX?

(what is TEX?). Wiadomości Matematyczne, 29(1):132–156, 1990. In Polish.

Billawala:STAN-CS-89-1256

[Bil89] Neenie Billawala. Metamarks:

Preliminary studies for a Pandora’s Box of shapes. Technical Report STAN-CS-89-1256, Stanford University, Department of Computer Science, May 1989.

Binding:MDM-4-19-22

[Bin93]  Angelika Binding.  Einsatz

von TEX beim Springer-Verlag. Mitteilungen der Deutschen Mathematiker-Vereinigung, 19: 19–22, 1993.

Bien:TEX85-133

[BK85] Janusz S. Bień and Hanna Kołodziejska. TEX for Riad Computers. In Lucarella [Luc85b], pages 133–140. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Bentley:EPODD-1-1-3

[BK88] J. L. Bentley and B. W. Kernighan. Tools for printing indexes. Electronic Publishing— Origination, Dissemination, and Design, 1(1):3–18, April 1988.

Bruggemann-Klein:EPODD-2-2-101

[BKW89] A. Brüggeman-Klein and D. Wood. Drawing trees nicely with TEX. Electronic Publishing—Origination, Dissemination, and Design, 2(2):101–115, July 1989.

Black:TDP90

[Bla90] Alison Black. Typefaces for desktop publishing; a user guide. Architecture Design and Technology Press, Ltd., 140 Kensington Church St., London W84 BN, UK, 1990. ISBN 1-85454-841-7. 106 pp. LCCN Z253.3.B53 1990.

Baecker:HFT89

[BM89] Ronald M. Baecker and Aaron Marcus. Human Factors and Typography for More Readable Programs. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-10745-7. xx + 348 pp. LCCN QA76.75 .B34 1990.

Borceux:LPT90

[Bor90] Francis Borceux. LaTEX—la perfection dans le traitement du texte. Editions Ciaoco, Artel, Bruxelles, Belgium, 1990. ISBN 2-87085-194-4.

Borde:TE92

[Bor92a] Arvind Borde. TEX by Example.

Academic Press, New York, NY, USA, 1992. ISBN 0-12-117650-9. xiv + 169 pp. LCCN Z253.4.T47 B67 1992. This book was designed by the author, with some assistance from the production department of Academic Press, Boston. The covers were done by the publisher. It was typeset by the author using Plain TEX plus specially written macros. The code for all macros used for the book are reproduced in it so that the book itself contains all the information—page layout specifications, fonts used, and so on— about its own typesetting. Over half the book consists of TEX input and output shown side-by-side, providing an explicit example of how a part of the book was typeset. The Academic Press logo on page iii, the catalogingin-publication data on page iv, and the diagrams on page 33 were pasted in place by the production department of Academic Press. The diagrams were drawn using LaTEX. Sample printouts of pages were made on various laser printers. When everything seemed satisfactory, the dvi files were copied onto IBM diskettes and were mailed to Academic Press, and from them to the American Mathematical Society. The final copy was produced by the A.M.S. using an APS micro-5 phototypesetter. All fonts used in the book are from the Computer Modern family. Parts of the covers of the book, inside and outside, were done in TEX by the author (the code for this is not included in the book). The outside cover was designed by Camille Pecoul of Academic Press and produced at Allison Associates in Boston. The font used for the title (on the covers) is Meridien. The background pattern was obtained by shooting a handmade marble paper as a black-and-white halftone and then overprinting with the Pantone Matching System (PMS). The placement of material on the cover was done by hand by Ms. Pecoul using a mechanical board. The drop shadow effect was done at Allison Associates.

Borkowski:LPS92

[Bor92b] Kazimierz M. Borkowski. LaTEX. Profesjonalny skład publikacji. Adam Marszałek, Toruń, Poland, 1992. In Polish.

Borde:MTE93

[Bor93] Arvind Borde. Mathematical TEX by Example. Academic Press, New York, NY, USA, 1993. ISBN 0-12-117645-2. xii + 352 pp. LCCN Z253.4.T47 B67 1993.

Boston:STT86

[Bos86] Bruce O. Boston. STET! Tricks of the Trade for Writers and Editors. Editorial Experts, Inc., Alexandria, VA, USA, 1986. ISBN 0-935012-07-9. 310 pp. LCCN PN162 .S741 1986. US$15.95.

Borde:UGT93

[BR93] Arvind Borde and Tomas Rokicki. A User’s Guide for TEXHelp; The On-Line TEX Handbook. Academic Press, New York, NY, USA, 1993. ISBN 0-12-117640-1. v + 28 pp. This small booklet documents the TEXHelp program for IBM PC and compatible microcomputers.

Braswell:IP89

[Bra89] Frank Merritt Braswell. Inside PostScript. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1989. ISBN 0-938151-10-X. LCCN QA76.73.P67 B73 1989. US$37.50.

Brooks:COMPUTER-24-6-7

[Bro91] Kenneth P. Brooks. Lilac: A two-view document editor. Computer, 24(6):7–19, June 1991.

Bryan:SAG88

[Bry88] Martin Bryan. SGML—An Author’s Guide to the Standard Generalized Markup Language. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-17535-5. xvii + 364 pp. LCCN QA76.73.S44 B79 1988.

Brandt:CAC89

[BU89] Josef Brandt and Ivar K. Ugi, editors. Computer Applications in Chemical Research and Education: A Collection of Papers Based on Lectures that were Presented at the VIIth International Conference on Computers in Chemical Research and Education at Garmisch-Partenkirchen, DE. Computeranwendung in der Chemie. Dr. Alfred Hüthing Verlag, Heidelberg, Germany, 1 edition, 1989. ISBN 3-7785-1804-6. xiii + 449 pp. LCCN QD39.3.E46 I56 1985. DM128.00.

Buerger:LES90

[Bue90] David J. Buerger. LaTEX for Engineers and Scientists. McGraw-Hill, New York, NY, USA, 1990. ISBN 0-07-008845-4. xvi + 198 pp. LCCN Z253.4.L38 B84 1990.

Butcher:CEC81

[But81] Judith Butcher. Copy editing: the Cambridge handbook. Cambridge University Press, New York, second edition, 1981. ISBN 0-521-23868-4, 0-521-25638-0. LCCN PN162 .B86 1981. UK£22.

Bzyl:TT93

[Bzy93]  Włodek Bzyl.  TEX. TEX??

broszura z konferencji członków GUSTu, Bachotek, 1993. In Polish.

GUTenberg:cahiers

[Cah ] Cahiers GUTenberg, 1988–. Journal of Groupe des Utilisateurs de TEX Francophones (group of French-speaking TEX Users).

Casti:NAMS-38-2-105

[Cas91] John L. Casti. VTEX typesetting package. Notices of the American Mathematical Society, 38(2):105– 109, February 1991.

Casti:NAMS-40-5-464

[Cas93] John L. Casti. Bitmap-free TEX.

Notices of the American Mathematical Society, 40(5):464–468, May/June 1993. A review of Y&Y’s DVIWindo and DVIP-SONE, and comments on outline fonts for TEX.

Canzii:TEX85-15

[CDMR85] G. Canzii, G. Degli Antoni, S. Mussi, and G. Rossi. SDDS: Scientific Document Delivery System. In Lucarella [Luc85b], pages 15–25. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Cerofolini:TEX85-191

[Cer85] Luigi Cerofolini. A Low Cost Laser Beam Printer Controller. In Lucarella [Luc85b], pages 191– 195. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Chen:UCB-TR-87-347

[CH87] Pehong Chen and Michael A.

Harrison. Automating index preparation. Technical Report 87/347, Computer Science Division, University of California, Berkeley, CA, USA, March 1987. This is an expanded version of [CH88].

Chen:SPE-19-9-897

[CH88] Pehong Chen and Michael A.

Harrison. Index preparation and processing. Software—Practice and Experience, 19(9):897–915, September 1988. The LaTEX text of this paper is included in the makeindex software distribution.

Cheswick:ATTCSTR-145

[Che90] Bill Cheswick. A permuted index for TEX and LaTEX. Computing Science Technical Report 145, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 1990. Also published by TUG as TEXniques No. 14.

Church:NAMS-39-8-842

[Chu92] P. T. Church. Using workshops to teach TEX. Notices of the American Mathematical Society, 39(8): 842, October 1992. AMS-TEX input for two workshop handouts may be obtained by anonymous ftp from syr.edu in the directory

                                        /math/papers.

Clark:TAU88

[Cla90] Malcolm Clark, editor. TEX: Applications, Uses, Methods. Proceedings, Third European TEX Conference, TEX88, Exeter, August 1988. Ellis Horwood, Chichester, UK, 1990. ISBN 0-13-912296-6. LCCN Z253.4.T47 C45 1990. US$69.95.

Clark:PTP92

[Cla92] Malcolm Clark. A Plain TEX Primer. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1992. ISBN 0-19-853784-0 (hardback), 0-19-853724-7 (softback). 481 pp. LCCN Z253.4.T47 C46 1992.

comp.text.tex

[com ] comp.text.tex, 1989–. This is an unmoderated Usenet discussion list about TEX.

Cornell:MI-8-2-64

[Cor86]  Gary Cornell.  Chiwriter:

Horstman software group. The Mathematical Intelligencer, 8(2): 64–65, 1986.

Craig:DT71

[Cra71] James Craig. Designing with Type. Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 1971. ISBN 0-8230-1320-0. 175 pp. LCCN 70-159564.

Dair:DT67

[Dai67] Carl Dair. Design with Type.

University of Toronto Press, Toronto, Ontario, Canada, 1967. ISBN 0-8020-1426-7 (hardcover), 0-8020-6519-8 (paperback). vi + 162 pp. LCCN Z246 D35 1967.

DANTE:tex-komodie

[DAN ] Die TEXnische Komödie, 1989– . Journal of DANTE (Deutschsprachige Anwendervereinigung TEX e.V.); Group of Germanspeaking TEX Users.

Darling:NAMS-40-2-150

[Dar93] R. W. R. Darling. Framemaker 3.0. Notices of the American Mathematical Society, 40(2):150– 152, February 1993.

Bruin:PL88

[dBvLV88] R. de Bruin, C. G. van der Laan, J. R. Luyten, and H. F.

Vogt. Publiceren met LaTEX. CWI Syllabus 19, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1988. ISBN 90-6196-357-5. In Dutch; copious examples.

Dreyfus:AFY89

[DE89] John Dreyfus and Knut Erichson, editors. ABC-XYZapf: Fifty Years in Alphabet Design: Professional and personal contributions selected for Hermann Zapf. The Wynkyn de Worde Society and Bund Deutscher Buchkünstler, London, UK and Offenbach, Germany, 1989. ISBN 3-922715-35-4. 251 pp. LCCN Z250.A2 Z358 1989.

Deach:SEYBOLD-6-7-21

[Dea92] Stephen Deach. Outline font hints and rasterization: A technology primer. The Seybold Report on Desktop Publishing, 6(7): 21–31, March 9 1992. This paper is followed by a dissent from Adobe (Terry O’Donnell), and a response from the author (p. 32). A sidebar (historical note, p.

25) contains a short description of and cites - as a forerunner of the hinting mechanism used in True-Type.

Desarmenien:STAN-CS-84-1013

[Dés84a] Jacques Désarménien. How to run TEX in French. Technical Report STAN-CS-84-1013, Stanford University, August 1984.

Desarmenien:SU-84

[Dés84b] Jacques Désarménien. La division par ordinateur des mots français avec le logiciel TEX. In How to run TEX in French, pages 19–42. Stanford University, Stanford, CA, USA, 1984.

Desarmenien:TEX85-41

[Dés85] Jacques Désarménien. The Use of TEX in French: Hyphenation and Typography. In Lucarella [Luc85b], pages 41–59. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Desarmenien:TSD86

[Dés86] Jacques Désarménien, editor.

TEX for Scientific Documentation. Second European Conference. Strasbourg, France, June 19–21, 1986. Proceedings, number 236 in Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 3-540-16807-9, 0-387-16807-9 (paperback). LCCN T 385 E97 1989.

Diller:LLL93

[Dil93] Antoni Diller. LaTEX Line by Line: Tips and Techniques for Document Processing. Wiley Professional Computing. Wiley, Chichester, UK, 1993. ISBN 0-471-93471-2 (softcover). xiii + 291 pp. LCCN Z253.4.L38 D55 1993. US$29.95. Optionally accompanied by disk with examples, ISBN 0-471-93797-5.

Duan:EPODD-2-1-47

[DM89] G. Duan and R. A. Morris. The importance of phase in the spectra of digital type. Electronic Publishing—Origination, Dissemination, and Design, 2(1): 47–60, April 1989.

DeMillo:SERC-TR-92-P

[DM91] Richard A. DeMillo and Aditya P. Mathur. On the use of software artifacts to evaluate the effectiveness of mutation analysis for detecting errors in production software. Technical Report SERC-TR-92-P, Software Engineering Research Center, Purdue University, W. Lafayette, IN 47907, USA, 1991.

Doob:NAMS-36-9-1203

[Doo89] Michael Doob. TEX and typesetting—an author’s view. Notices of the American Mathematical Society, 36(9):1203–1204, November 1989.

Doob:NAMS-37-3-270

[Doo90] Michael Doob. TEX and the single CPU. Notices of the American Mathematical Society, 37(3): 270–273, March 1990.

Doob:TS193

[Doo93] Michael Doob. TEX: Starting from 1. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-56441-1. 114 pp. LCCN Z253.4.T47 D66 1993. US$25.00.

Dyl:KST94

[Dyl94]  Dariusz Dyl.  Komputerowy

skład tekstu w systemie LaTEX. Przewodnik po komendach i środowiskach wraz z przykładami. Adam Marszałek, Toruń, Poland,

1. edition, 1994. składano w systemie LaTEX (styl book11pt) przy użyciu programu VectorTEX ver. 2.24.

Eickel:LLS89

[Eic90] Jürgen Eickel. Logical and layout structures of documents. In Nadrchal [Nad90], pages 201– 208. Invited paper.

Eijkhout:TTT92

[Eij92] Victor Eijkhout. TEX by Topic, A TEXnician’s Reference. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56882-9. vii + 307 pp. LCCN Z253.4.T47 E38 1992.

English:MF-design

[Eng87]  Paul M. English. Using  -

for original font design. August 1987.

Eijkhout:TWIOscoop-8-2-44

[EP90] Victor Eijkhout and Nico Poppelier. Wat is TEX. TWIOscoop, 8 (2):44–48, 1990.

EPODD

[EPO ] Electronic publishing—origination, dissemination, and design. Wiley, 1988–. ISSN 0894-3982. This journal accepts papers in TEX form.

Fateman:SIGSAM-21-4-1

[Fat87] Richard J. Fateman. TEX output from MACSYMA-like systems. SIGSAM Bulletin, 21(4): 1–5, November 1987.

CT89

[Féd89] Fédération C. G. C. de la Communication. Code Typographique —Choix de règles à l’usage des auteurs et professionels du livre, seizième edition, 1989. This book is a compendium of rules for French typography, and for the typesetting of several foreign languages in French texts. It includes tables of diacritics for most European languages, and hyphenation rules for English, German, Greek, Italian, Latin, Portuguese, and Spanish.

Fuchs:STAN-CS-82-901

[FK82] David R. Fuchs and Donald E. Knuth. Optimal Font Caching. Technical Report STAN-CS-82-901, Stanford University, 1982.

Felici:DPS87

[FN87] James Felici and Ted Nace. Desktop Publishing Skills. Addison-Wesley, Reading, MA, USA, August 1987. ISBN 0-201-11537-9. vii + 180 pp. LCCN Z286.D47 F44 1987. UK£20.

Foata:LTI-84

[FPR84] Dominique Foata, Jean-Jacques Pansiot, and Yves Roy. STRATEC —un logiciel traitement de textes mathématiques en amont de TEX (ver.0.93). Laboratoire de Typographie Informatique, Université Louis Pasteur, Strasbourg, France, 1984.

Foata:TEX85-105

[FPR85] Dominique Foata, Jean-Jacques Pansiot, and Yves Roy. STRATEC and a Rationalized Keyboard for Inputting TEX. In Lucarella [Luc85b], pages 105–116. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Furuta:COMP-SURV-14-3-417

[FSS82a] Richard Furuta, Jeffrey Scofield, and Alan Shaw. Document formatting systems: survey, concepts, and issues. ACM Computing Surveys, 14(3):417–472, September 1982.

Furuta:DFS82

[FSS82b] Richard Furuta, Jeffrey Scofield, and Alan Shaw. Document formatting systems: survey, concepts, and issues, chapter ??, pages 133–220. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1982. ISBN ?? LCCN ??

Fujita:COMPCHEM-18-2-109

[Fuj94] Shinsaku Fujita. Typesetting structural formulae with the text formatter TEX/LaTEX. Computers in Chemistry, 18(2):109–116, 1994.

OrtegonGallego:LPP92

[Gal92] Francisco Ortegón Gallego. LaTEX, primeros pasos. Manuales de Informática Masson. Masson, S.A., Barcelona, Spain, 1992. ISBN 84-311-0649-2. xix + 168 pp.

Ghosh:STAN-CS-83-966

[GB83] Pijush K. Ghosh and Charles A. Bigelow. A Formal Approach to Lettershape Description for Type Design. Technical Report STAN-CS-83-966, Stanford University, 1983.

Griffee:IBMSYSJ-27-2-206

[GC88] Alan W. Griffee and Cherrelyn A. Casey. An introduction to typographic fonts and digital font resources. IBM Systems Journal, 27(2):206, 1988.

Gehani:DFT87

[Geh87] Narain Gehani. Document Formatting and Typesetting on the UNIX System. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, second edition, 1987. ISBN 0-13-938325-5. xix + 310 pp. LCCN QA76.73.A35 G437 1987.

Guoan:STAN-CS-82-914

[GH82] Gu Guoan and John Hobby. Using String Matching to Compress Chinese Characters. Report No. STAN-CS-82-914, Stanford University, Department of Computer Science, 1982.

Ghosh:STAN-CS-83-965

[Gho83] Pijush K. Ghosh. An Approach to Type Design and Text Composition in Indian Scripts (with a preface by donald E. knuth). Technical Report STAN-CS-83-965, Stanford University, 1983.

Gilchrist:NAMS-36-9-1199

[Gil89] Martin Gilchrist. TEX and typesetting—A publisher’s view. Notices of the American Mathematical Society, 36(9):1199–1201, November 1989.

Goosens:LC94

[GMS94] Michel Goossens, Frank Mittelbach, and Alexander Samarin. The LaTEX Companion. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-54199-8. xxi + 528 pp. LCCN Z253.4.L38 G66 1994. US$34.25.

Gourlay:SPE-29-5-388

[Gou86] John S. Gourlay. A language for music printing. Software—Practice and Experience, 29(5):388– 401, May 1986.

Gonczarowski:EPODD-2-3-157

[GP89] Jakob Gonczarowski and On G.

Paradise. InScript—a C-like preprocessor for PostScript. Electronic Publishing—Origination, Dissemination, and Design, 2(3): 157–167, October 1989.

Garvey:NEM82

[GR82] Daniel E. Garvey and William L.

Rivers. Newswriting for the Electronic Media: Principles, Examples, Applications. Wadsworth Publishing Company, Belmont, CA, USA, 1982. ISBN 0-534-01069-5. xiii + 250 pp. LCCN PN4784.B75 .G3. Exercises, appendices, index, pp. 250.

Gratzer:MT92

[Grä93a] George Grätzer. Math into TEX: A Simplified Introduction Using AMS-LaTEX. Springer-Verlag and Birkhäuser, Berlin, Germany / Heidelberg, Germany / London, UK / etc. and Basel, Switzerland, 1992 (or 1993??). ISBN 0-8176-3637-4, 3-7643-3637-4. 187 (or xxix + 294)?? pp. LCCN Z253.4.A65G7 1993. US$34, SFr 52, DM 57.

Gratzer:NAMS-40-2-148

[Grä93b] George Grätzer. AMS-LaTEX. Notices of the American Mathematical Society, 40(2):148–150, February 1993.

Gratzer:NAMS-40-7-834

[Grä93c] George Grätzer. Advances in TEX implementations. I. postscript fonts. Notices of the American Mathematical Society, 40(7):834–838, September 1993.

Guntermann:TEX85-197

[Gun85] Klaus Guntermann. GENDRIV— a driver for low cost devices using built-in fonts. In Lucarella [Luc85b], pages 197–204. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Gurari:TLD94

[Gur94a] Eitan M. Gurari. TEX and LaTEX: Drawing and Literate Programming. McGraw-Hill, New York, NY, USA, 1994. ISBN 0-07-025208-4 (book), 0-07-911616-7 (book + diskette). xiv + 310 pp. LCCN Z253.4.T47G87 1994. US$34.95.

Gurari:WT94

[Gur94b] Eitan M. Gurari. Writing with TEX. McGraw-Hill, New York, NY, USA, 1994. ISBN 0-07-025207-6. xiv + 249 pp. LCCN Z253.4.T47 G87 1994. US$29.95.

Hegron:IRISA-244

[H´84] Gérard Hégron. Algorithmes de génération de caractères. Publication interne 244, IRISA, Rennes, France, 1984.

Hahn:LE91

[Hah91] Jane Hahn. LaTEX for Everyone—A Reference Guide and Tutorial for Typesetting Documents Using a Computer. Personal TEX Inc., 12 Madrona Street, Mill Valley, CA 94941, USA, 1991. ISBN 0-9631044-0-3. xi + 346 pp. LCCN Z253.4.L38 H34 1991. US$19.95.

Hahn:LE93

[Hah93] Jane Hahn. LaTEX for Everyone—A Reference Guide and Tutorial for Typesetting Documents Using a Computer. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1993. ISBN 0-13-605908-2. xi + 346 pp. LCCN Z253.4.L38 H35 1993. US$19.95.

Harrower:NDH91

[Har91a] Tim Harrower. The Newspaper Designer’s Handbook. William C. Brown Company Publishers, Dubuque, IA, USA, 1991.

Hart:HRC91

[Har91b] Horace Hart. Hart’s Rules; For Compositors and Readers at the University Press, Oxford. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 39th edition, 1991. ISBN 0-19-212983-X. 182 pp. LCCN Z254 .H37 1983. UK£8.

Hayward:SA77

[Hay77] Stan Hayward. Scriptwriting for Animation. Hastings House, Publishers, New York, NY, USA, 1977. ISBN 0-8038-6741-7. 160 pp.

Hayashi:SPE-17-1-17

[Hay87] T. Hayashi. Transporting TEX into the HITAC VOS3 environment. Software—Practice and Experience, 17(1):17–28, January 1987.

Heinemann:XPS88

[HC88] Sue Heinemann and Virginia Croft, editors. Xerox Publishing Standards: A Manual of Style and Design. Watson-Gup-till Publications, 1515 Broadway, New York, NY 10036, USA, 1988. ISBN 0-8230-5964-2. LCCN Z286.L68 X47 1988. US$35.00.

Heck:DPAS92

[Hec92] André Heck, editor. Desktop Publishing in Astronomy & Space Sciences. Strasbourg Astronomical Observatory, [1–3] October, 1991. World Scientific, Singapore, 1992. ISBN 981-02-0915-0. LCCN QB14.2 .D47 1992. In addition to the papers presented at the colloquium, each paper is followed by a summary of the discussion about that paper.

Hershey:TR-2339

[Her69] Allen V. Hershey. FORTRAN IV programming for cartography and typography. Technical Report TR-2339, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, September 1969.

Hershey:TN-K-27-70

[Her70] Allen V. Hershey. Preparation of reports with the FORTRAN typographic system. Technical Report TN-K/27-70, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, September 1970.

Hershey:CGIP-1-373

[Her72] Allen V. Hershey. A computer system for scientific typography. Computer Graphics and Image Processing, 1:373–385, 1972.

Hershey:ACT81

[Her81] Allen V. Hershey. Advanced computer typography. Technical Report NPS012-81-005, U. S. Naval Postgraduate School, Monterey, CA 93940, USA, December 1981.

Hobby:STAN-CS-83-974

[HG83] John Hobby and Gu Guoan. A Chinese Meta-Font. Report No. STAN-CS-83-974, Stanford University, Department of Computer Science, 1983.

Higham:HWM93

[Hig93] Nicholas J. Higham. Handbook of Writing for the Mathematical Sciences. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, June 1993. ISBN 0-89871-314-5. xii + 241 pp. LCCN QA42.H54 1993. US$21.50. [From the publisher]: Having trouble with your latest math paper? Giving a presentation that you just can’t pull together? Struggling with your thesis or trying to get your first article published in a technical journal? Handbook of Writing for the Mathematical Sciences is the book for you! This handy volume provides information on virtually every issue you will face when writing a technical paper or talk, from choosing the right journal to handling your references. You’ll also get an overview of the entire publication process– invaluable for anyone hoping to publish in a technical journal. To write a truly impressive paper, you’ll need to understand the anatomy of a research paper and the steps involved in revising a draft. This book offers discussions of these fundamental topics, along with illustrative and provocative examples. Also included are chapters on standard English usage, using computers for writing and research, and writing technical material when English is a foreign language. This handbook provides much-needed advice on handling the basic ingredients of a research paper, like definitions, theorems, examples, and equations. In addition, appendices provide essential reference material, including summaries of LaTEX symbols and Emacs commands, addresses of mathematical societies, and a list of papers that have won expository writing prizes. This book is ideal for graduate students and teachers. Among its special features:

• Detail and extensive use of examples make it an excellent teaching tool

• Discusses TEX and other software tools for preparing publications

• Thorough treatment, with examples, of how to write slides (transparencies) for a mathematical talk.

• Comprehensive index and extensive bibliography make it an excellent reference

• Readable for everyone in the field–from undergraduates to seasoned professionals

Contents Preface; Chapter 1: General Principles; Chapter 2: Writer’s Tools and Recommended Reading; Chapter 3: Mathematical Writing; Chapter 4: English Usage; Chapter 5: When English is a Foreign Language; Chapter 6: Writing a Paper; Chapter 7: Revising a Draft; Chapter 8: Publishing a Paper; Chapter 9: Writing a Talk; Chapter 10: Computer Aids for Writing and Research; Appendix A: The Greek Alphabet; Appendix B: Summary of TEX and LaTEX Symbols; Appendix C: GNU Emacs–The Sixty+ Most Useful Commands; Appendix D: Mathematical Organizations in the UK and USA; Appendix E: Winners of Prizes for Expository Writing; Appendix F: Glossary; Bibliography; Index. About the Author: Nicholas J. Higham is a Reader in Mathematics at the University of Manchester, UK. He is the author of more than 40 publications and is a member of the editorial board of the SIAM Journal on Matrix Analysis and Applications.

Hohti:SIGAPL-18-3-13

[Hoh88] Aarno Hohti. Typesetting APL with TEX. ACM SIGAPL APL Quote Quad, 18(3):13–16, March 1988.

Holzgang:UPP87

[Hol87] David A. Holzgang. Understanding PostScript Programming. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1987. ISBN 0-89588-396-1. xxxii + 459 pp. LCCN QA76.73.P67 H65 1987.

Holzgang:PPR89

[Hol89] David A. Holzgang. PostScript Programmer’s Reference Guide. Scott, Foresman and Company, Glenview, IL, USA, 1989. ISBN 0-673-38574-4. x + 486 pp. LCCN QA76.73.P67 H64 1989.

Huggins:STAN-CS-89-1251

[Hug88] K. Cleo R. Huggins. Egyptian hieroglyphs for modern printing devices. Technical Report STAN-CS-89-1251, Stanford University, June 1988.

Hurlburt:GMS78

[Hur78] Allen Hurlburt. The grid: A modular system for the design and production of newspapers, magazines, and books. Van Nostrand Reinhold, New York, NY, USA, 1978. ISBN 0-442-23598-4. 96 pp. LCCN Z246 .H82. US$15.00.

Ibrus:VET93

[Ibr92]  Hans Ibrus.  Väike Eesti

TEXiraamat. a/S “KompuT” kirjastus, Tartumaa, Estonia, 1992. 144 pp. In Estonian. Appendix A is a translation into Estonian of [Sam83].

Ikarus:83

[Ika83] IKARUS for Typefaces in Digital Form. URW Unternehmensberatung Karow Rubow Weber GMBH, Hamburg, Germany, 1983.

Incerpi:INRIA-71

[IM86a] Janet Incerpi and Francis Montagnac. Getting Faster TEX Output. Using the DVI Library. Rapports Techniques No. 71, INRIA/IRISA, Rennes, France, July 1986.

Incerpi:GFT86

[IM86b] Janet Incerpi and Francis Montagnac. Using the DVI Library. In Getting Faster TEX Output. INRIA/IRISA, Rennes, France, 1986.

Jackowski:MZP93

[Jac93]  Bogusław Jackowski.  Meta-

font: zastosowania praktyczne i niepraktyczne. broszura z konferencji członków GUSTu, Bachotek, 1993. In Polish.

Jackowski:WTP88

[JHR88] Bogusław Jackowski, Tomasz Hołdys, and Marek Ryćko. With TEX to the Poles: a few parochial remarks on using TEX by the non-English. Anima Ltd, Gdańsk, 1988.

Johnstone:LC92

[Joh92] Adrian Johnstone. LaTEX, Concisely. Ellis Horwood, Chichester, UK, 1992. ISBN 0-13-524539-7. xviii + 170 pp. LCCN Z253.4.L38 J64 1992. US$30.00.

Karow:DFT87

[Kar87] Peter Karow. Digital Formats for Typefaces. URW (Unternehmensberatung Rubow Weber) Verlag, Hamburg, Germany, 1987. ISBN 3-926515-01-5. 400 pp. LCCN Z253.3 .K371 1987.

Karow:DS92

[Kar92] Peter Karow. Digitale Schriften, Darstellung und Formate, Geleitwort von Hermann Zapf (Digital Fonts, Representation and Formats, Forward by Hermann Zapf). Springer-Verlag and URW (Unternehmensberatung Rubow Weber) Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc. and Hamburg, Germany, 1992. xiii + 457, with 230 illustrations pp.

Karow:F94

[Kar94] Peter Karow. Fonttechnology.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-57223-6. 450 pp. US$56.00.

Katzenbeisser:ET93

[Kat93] Stefan Katzenbeisser. Einführung in TEX. R. Oldenbourg Verlag, Munich, Germany and Vienna, Austria, 1993. ISBN 3-486-22682-7. 108 pp.

Kopka:GLD93

[KD93] Helmut Kopka and Patrick W.

Daly. A Guide to LaTEX: Document Preparation for Beginners and Advanced Users. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-56889-6. US$34.50.

Keller:TEX85-1

[Kel85] Arthur M. Keller. Typesetting by Authors. In Lucarella [Luc85b], pages 1–14. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Knuth:CSS93

[KL93] Donald E. Knuth and Silvio Levy. The CWEB System of Structured Documentation, Version 3.0. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-57569-8. 226 pp. LCCN QA76.9.D3 K6 1993.

Knuth:MW89

[KLR89] Donald E. Knuth, Tracy Larrabee, and Paul M. Roberts. Mathematical Writing. Mathematical Association of America Notes Number 14, 1989.

Knuth:STAN-CS-78-648

[Knu78a] Donald E. Knuth. Mathematical typography. Technical Report STAN-CS-78-648, Stanford University, February 1978. Also published as [Knu79b].

Knuth:STAN-CS-78-675

[Knu78b] Donald E. Knuth. TAU EP-SILON CHI. A System for Technical Text. Report No. STAN-CS-78-675, Stanford University, Department of Computer Science, 1978.

Knuth:TMN79

[Knu79a] Donald E. Knuth. TEX and —New Directions in Typesetting. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1979. ISBN 0-932376-02-9. xi + 201 + 105 pp. LCCN Z253.3 .K58 1979.

Knuth:BAMSN-1-337

[Knu79b] Donald E. Knuth. Mathematical typography. Bulletin of the American Mathematical Society (new series), 1:337–372, March 1979. Also published as a Stanford CS Report, STAN-CS-78-

Knuth:STAN-CS-79-762

[Knu79c] Donald E. Knuth. Metafont—A System for Alphabet Design. Report No. STAN-CS-79-762, Stanford University, Department of Computer Science, 1979.

Knuth:MI-2-3-114

[Knu80a] Donald E. Knuth. The letter S.

The Mathematical Intelligencer, 2(3):114–112, 1979, 1980. Also published as [Knu80c].

Knuth:STAN-CS-80-780

[Knu80b] Donald E. Knuth. The Computer Modern Family of Typefaces. Report No. STAN-CS-80-780, Stanford University, Department of Computer Science, 1980.

Knuth:STAN-CS-80-795

[Knu80c] Donald E. Knuth. The letter S. Technical Report STAN-CS-80-795, Stanford University, April 1980. Also published as [Knu80a].

Knuth:STAN-CS-81-886

[Knu81] Donald E. Knuth. The concept of a meta-font. Technical Report STAN-CS-81-886, Stanford University, October 1981.

Knuth:tex82

[Knu82a] Donald E. Knuth. TEX82. Stanford University, Stanford, CA, USA, 1982.

Knuth:VL-16-1-3

[Knu82b] Donald E. Knuth. The Concept of a Meta-Font. Visible Language, 16(1):3–27, winter 1982.

Knuth:TFT82

[Knu82c] Donald E. Knuth. A test file for TEX. Stanford University, Stanford, CA, USA, 1982.

Knuth:texware82

[Knu82d] Donald E. Knuth. TEXware (version 1). Stanford University, Stanford, CA, USA, 1982.

Knuth:web82

[Knu82e] Donald E. Knuth. The WEB System of Structured Documentation—Version 1. Stanford University, Stanford, CA, USA, 1982.

Knuth:STAN-CS-83-978

[Knu83a] Donald E. Knuth. Lessons Learned from . Report No. STAN-CS-83-978, Stanford University, Department of Computer Science, 1983.

Knuth:STAN-CS-83-981

[Knu83b] Donald E. Knuth. Literate programming. Report No. STAN-CS-83-981, Stanford University, Department of Computer Science, 1983.

Knuth:STAN-CS-83-980

[Knu83c] Donald E. Knuth. The WEB System of Structured Documentation—ver.2.3. Report No. STAN-CS-83-980, Stanford University, Department of Computer Science, 1983.

Knuth:TB84

[Knu84a] Donald E. Knuth. The TEXbook. Addison-Wesley, Reading, MA, USA, 1984. ISBN 0-201-13448-9 (paperback), 0-201-13447-0 (hardcover). ix + 483 pp. LCCN Z253.4.T47 K58 1984. US$15.95 (paperback), US$32.95 (hardcover).

Knuth:CJ-2-27-97

[Knu84b] Donald E. Knuth. Literate programming. The Computer Journal, 27(2):97–111, May 1984.

Knuth:STAN-CS-84-1027

[Knu84c] Donald E. Knuth. A torture test for TEX. Report No. STAN-CS-84-1027, Stanford University, Department of Computer Science, 1984.

Knuth:INRIA-119-85

[Knu85] Donald E. Knuth. Le concept de Metafonte. In André [And85], pages 119–132. Przedruk z “Communication et langages” no.55 (1983) pp.40–53.

Knuth:STAN-CS-86-1095

[Knu86a] Donald E. Knuth. A Torture Test for . Report No. STAN-CS-86-1095, Stanford University, Department of Computer Science, 1986.

Knuth:STAN-CS-86-1097

[Knu86b] Donald E. Knuth. TEXware.

Report No. STAN-CS-86-1097, Stanford University, Department of Computer Science, April 1986.

Knuth:ct-a

[Knu86c] Donald E. Knuth. The TEXbook, volume A of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13447-0. ix + 483 pp. LCCN Z253.4.T47 K58 1986. US$32.95.

Knuth:ct-b

[Knu86d] Donald E. Knuth. TEX: The Program, volume B of Computers and Typesetting. Addison-Wes-ley, Reading, MA, USA, 1986. ISBN 0-201-13437-3. xv + 594 pp. LCCN Z253.4.T47 K578 1986. US$32.95.

Knuth:ct-c

[Knu86e] Donald E. Knuth. The - book, volume C of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13445-4 (hardcover), 0-201-13444-6 (paperback). xi + 361 pp. LCCN Z250.8.M46 K58 1986. US$32.95 (hardcover), US$16.95 (paperback).

Knuth:ct-d

[Knu86f] Donald E. Knuth. : The Program, volume D of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13438-1. xv + 560 pp. LCCN Z250.8.M46 K578 1986. US$32.95.

Knuth:ct-e

[Knu86g] Donald E. Knuth. Computer Modern Typefaces, volume E of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13446-2. xv + 588 pp. LCCN Z250.8.M46 K574 1986.

Knuth:TOGS-6-4-245

[Knu87]  Donald E. Knuth.  Digital

halftones by dot diffusion. ACM Transactions on Graphics, 6(4): 245–273, October 1987.

Knuth:STAN-CS-88-1223

[Knu88] Donald E. Knuth. The errors of TEX. Technical Report STAN-CS-88-1223, Stanford University, Department of Computer Science, September 1988.

Knuth:TB89

[Knu89a] Donald E. Knuth. The TEXbook.

ASCII Corporation, Tokyo, Japan, 1989. ISBN 4-7561-0010-4. Translation of the TEXbook into Japanese, with addition of Appendix K, and new drawings by Duane Bibby.

Knuth:TEXHAX-89-98

[Knu89b] Donald E. Knuth. Calling all grand wizards. TEXhax, 89(98), November 1989.

Knuth:SPE-19-7-607

[Knu89c] Donald E. Knuth. The errors of TEX. Software—Practice and Experience, 19(7):607–681, July 1989. This is an updated version of [Knu88].

Knuth:TB11-1-13-23

[Knu90a] Donald Knuth. Virtual Fonts:

More Fun for Grand Wizards. TUGboat, 11(1):13–23, April 1990.

Knuth:TEXHAX-90-11

[Knu90b] Donald E. Knuth. Virtual fonts:

More fun for Grand Wizards.

TEXhax, 90(11 and 12), January 1990.

Knuth:LP92

[Knu92] Donald E. Knuth. Literate Programming. CSLI Lecture Notes Number 27. Stanford University Center for the Study of Language and Information, Stanford, CA, USA, 1992. ISBN 0-937073-80-6 (paperback), 0-937073-81-4 (hardcover). xiii + 368 pp. LCCN QA76.6 .K644 1992. US$24.95 (paperback), US$59.95 (hardcover).

Kolodziejska:dzielenie

[Koł87] Hanna Kołodziejska. Dzielenie wyrazów polskich w systemie TEX. Technical Report 165, Sprawozdania Instytutu Informatyki Uniwersytetu Warszawskiego, 1987.

Kolodziejska:MIKROLAN-1-12

[Koł88a]  Hanna Kołodziejska.  TEX.

Mikrolan, (1):12, 1988.

Kolodziejska:CG-0-3

[Koł88b] Hanna Kołodziejska. Le traitement des textes polonais avec le logiciel TEX. Cahiers GUTenberg, (0):3–10, April 1988.

Kopka:LE88

[Kop88] Helmut Kopka. LaTEX—Eine Einführung. Addison-Wesley Verlag, Bonn, Germany, second edition, 1988. ISBN 3-89319-199-2. xvi + 340 pp.

Kopka:LE90

[Kop90] Helmut Kopka. LaTEX—Erweiterungsmöglichkeiten. Addison-Wesley Verlag, Bonn, Germany, 1990. ISBN 3-89319-287-5.

Kopka:LEE91

[Kop91a] Helmut Kopka. LaTEX—Erweiterungsmöglichkeiten mit einer Einfürung in . Addison-Wesley Verlag, Bonn, Germany, second edition, 1991. ISBN 3-89319-356-1. 552 pp.

Kopka:LE91

[Kop91b] Helmut Kopka. LaTEX—Eine Einführung. Addison-Wesley Verlag, Bonn, Germany, third edition, 1991. ISBN 3-89319-338-3. 375 pp.

Kopka:LE93

[Kop93] Helmut Kopka. LaTEX—Eine Einführung. Addison-Wesley Verlag, Bonn, Germany, fourth edition, 1993. ISBN 3-89319-434-7. xviii + 461 pp.

Kopka:LE94a

[Kop94] Helmut Kopka. LaTEX Band 1: Einführung. Addison-Wes-ley Verlag, Bonn, Germany, 1994. ISBN 3-89319-664-1. xix + 428 pp. Replaces the old series of two volumes.

Kopka:LE94b

[Kopara] Helmut Kopka. LaTEX Band 2:

Ergänzungen. Addison-Wesley Verlag, Bonn, Germany, 1994 (to appear).

Kopka:LE94c

[Koparb] Helmut Kopka. LaTEX Band 3: Erweiterungen—mit einer Einfürung in . Addison-Wesley Verlag, Bonn, Germany, 1994 (to appear).

Knuth:SPE-11-1119

[KP81] Donald E. Knuth and Michael F. Plass. Breaking Paragraphs into Lines. Software—Practice and Experience, 11(11):1119– 1184, November 1981.

Kronert:IIS89

[Krö90] Günther Krönert. Importance of the ISO standard 8613 for document interchange. Basic standard, functional standards, initial experience, benefits. In Nadrchal [Nad90], pages 209–218. Invited paper.

Krieger:IT90

[KS90] Jost Krieger and Norbert Schwarz. Introduction to TEX. Addison-Wesley Europe, Amsterdam, The Netherlands, 1990. ISBN 0-201-51141-X. English translation of [Sch88b]. See also the Dutch translation, [Sch90].

Kernighan:COMPSYS-2-2-103

[KW89] Brian W. Kernighan and Christopher J. Van Wyk. Page makeup by postprocessing text formatter output. Computing Systems, 2 (2):103–132, 1989.

Knuth:TEXHAX-90-13

[KW90] Donald E. Knuth and Joe Weening. New TEX/ sources available on Stanford’s master archive. TEXhax, 90(13), January 1990.

Knuth:SP-20-3-131

[KZ89] Donald E. Knuth and Hermann Zapf. AMS euler—A new typeface for mathematics. Scholarly Publishing, 20(3):131–157, April 1989.

Labuz:HTW84

[Lab84] Ronald Labuz. How to Typeset from a Word Processor: A Practical Manual for Type Buyers and Typesetters. R. R. Bowker Company, New York, NY, USA, 1984. ISBN 0-8352-1899-6. xii + 218 pp. LCCN Z253.3.L33 1984. US$24.95.

Lamport:LDP86

[Lam85]  Leslie Lamport.  LaTEX—A

Document Preparation System —User’s Guide and Reference Manual. Addison-Wesley, Reading, MA, USA, 1985. ISBN 0-201-15790-X. xiv + 242 pp. LCCN Z253.4.L38 L35 1986. US$26.95.

Lao:WM-27-1

[Lao86] Marek J. Lao. Sztuka TEXowania.

Wiadomości Matematyczne, 27 (1):81–87, 1986. Roczniki Polskiego Towarzystwa Matematycznego, Seria II. In Polish.

Larsen:LD89

[Lar89] Steen Larsen. LaTEX på dansk.

UNI•C, Danmarks EDB-Center for Forskning og Uddannelse, Copenhagen, Denmark, 1989. ISBN 87-7252-089-2.

Lawson:AT90

[Law90] Alexander Lawson. Anatomy of a Typeface. Hamish Hamilton, London, 1990. ISBN 0-241-13267-3. 428 pp.

Lee:B79

[Lee79]  Marshall Lee.  Bookmaking:

the Illustrated Guide to Design/Production/Editing. R. R. Bowker Company, New York, NY, USA, second edition, 1979. ISBN 0-8352-1097-9. 485 pp. LCCN Z116.A2 .L44 1979. US$25.00.

Leitch:TEX85-171

[Lei85]  S. Leitch. Implementing  -

on an ICL Perq. In Lucarella [Luc85b], pages 171– 182. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Lesk:COMPSYS-1-3-207

[Les88] Michael Lesk. GRAB—inverted indexes with low storage overhead. Computing Systems, 1(3): 207–220, 1988.

Levine:CPC-58-1-181

[Lev90] Michael J. S. Levine. A LaTEX graphics routine for drawing feynman diagrams. Computer Physics Communications, 58(1 and 2):181–198, February/March 1990. Available in the CPC li-

                                        brary, and from listral@ib.rl.

ac.uk with an e-mail request GET

                                        CPC INTRO CPCINDEX.

Li:TEX85-161

[Li85] Jiarong Li. Generation of Some Chinese Characters with Metafont. In Lucarella [Luc85b], pages 161–170. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Liang:STAN-CS-83-977

[Lia83] Franklin Mark Liang. Word hy-phen-a-tion by com-pu-ter. Technical Report STAN-CS-83-977, Stanford University, August 1983.

Leitch:TEX85-155

[LS85] S. Leitch and F. J. Smith. Generalised Algorithm for Drawing Non-Parametric Splines. In Lucarella [Luc85b], pages 155–161. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Licha:A92

[LU92] Michaela Lichá and Oldřich Ulrych. AMS-TEX verze 2.1. Czechoslovak TEX Users Group, CSTUG MÚ UK, Sokolovská 83, 186 00 Praha 8, The Czech Republic, 1992. 80 pp. This publication provides a short (but complete) tour through AMS-TEX 2.1 together with amsppt. sty 2.1. Almost all features are demonstrated with a lot of examples. The book is organized in such a way that it is useful for novices (basic textbook), and for advanced users (quick reference guide) as well. Contains a table of all symbols.

Lubkin:CS-642

[Lub90] Saul Lubkin. Porting sophisticated programs to your Unix environment for free. Computer Shopper, pages 642–644, March

Lucarella:TEX85-93

[Luc85a] Dario Lucarella. TEX Formulae Dictionary. In TEXSCIDOC85 [Luc85b], pages 93–103. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Lucarella:TSD85

[Luc85b] Dario Lucarella, editor. Proceedings of the First European Conference on TEX for Scientific Documentation, 16–17 May 1985, Como, Italy. Addison-Wes-ley, Reading, MA, USA, 1985. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Luna:UTD92

[Lun92] Paul Felix John Luna. Understanding Type for Desktop Publishing. Blueprint, London, UK, 1992. ISBN 0-948905-76-X. 136 pp. Highly regarded English typographer: designed the Oxford English Dictionary and the New English Bible.

Linton:IEEE-COMPUTER-22-2-8

[LVC89] Mark A. Linton, John M. Vissides, and Paul R. Calder. Composing user interfaces with interviews. IEEE Computer, 22(2):8– 22, February 1989. From Computing Reviews: “InterViews objects are combined through a model similar to the boxes and glues in Knuth’s TEX, which allows interface objects to be combined with little concern for undesired interactions.”.

MacKay:TASxx

[Mac] Pierre A. MacKay. TEX for Arabic script. In ???, pages 391–400. ??, ??, ??

Manson:SIGDOC-13-1-3

[Man87]  David Manson.  Using TEX

for electronic publishing at I. P. Sharp Associates. ACM SIGDOC Asterisk*, 13(1):3–7, March 1987.

Morris:NIPT-344

[MBHL91] Robert A. Morris, Karl Berry, Kathryn A. Hargreaves, and Dimitrios Liarokapis. How typeface variation and typographic variation affect readability at small sizes. In Pietrowski [Pie91], pages 344–352.

McLean:THM80

[McL80] Ruari McLean. The Thames and Hudson Manual of Typography. Thames and Hudson, London, UK, 1980. ISBN 0-500-68022-1. 216 pp.

MF:MFC86

[MF86a] The MF Chel Family of Fonts.

The Metafoundry, OCLC Online Computer Library Center, Inc., Dublin, OH, USA, 1986.

MF:MFR86

[MF86b] The MF Roman Family of Fonts.

The Metafoundry, OCLC Online Computer Library Center, Inc., Dublin, OH, USA, 1986.

MF:MFD87

[MF87a] The MF Decorative Family of Fonts. The Metafoundry, OCLC Online Computer Library Center, Inc., Dublin, OH, USA, 1987.

MF:MFS87

[MF87b] The MF Slavic Families of Fonts. The Metafoundry, OCLC Online Computer Library Center, Inc., Dublin, OH, USA, 1987.

Miller:PROTEXT-I

[Mil84] John J. H. Miller, editor. PRO-TEXT I: Proceedings of the First International Conference on Text Processing Systems held in Dublin from 24th to 26th October, 1984. Boole Press, Dublin, Ireland, 1984. ISBN 0-906783-41-0 (hardcover), 0-906783-42-9 (paper). LCCN QA76.9.T48 I591 1984.

Miller:PROTEXT-II

[Mil85] John J. H. Miller, editor. PRO-TEXT II: Proceedings of the Second International Conference on Text Processing Systems held in Dublin from 23rd to 25th October, 1985. Boole Press, Dublin, Ireland, 1985. ISBN 0-906783-50-X (hardcover), 0-906783-40-2 (paperback). LCCN QA76.9.T48 I561 1985.

Miller:PROTEXT-III

[Mil86a] John J. H. Miller, editor. PRO-TEXT III: Proceedings of the Third International Conference on Text Processing Systems held in Dublin from 22nd to 24th October, 1986. Boole Press, Dublin, Ireland, 1986. ISBN 0-906783-57-7 (hardcover), 0-906783-58-5 (paperback). LCCN QA76.9.T48 I561 1986.

Milne:MI-8-1-66

[Mil86b] J. S. Milne. Review of two scientific word-processing programs: T^{3} version 2.02 by TCI Software Research Inc., and Volkswriter Scientific version 1.0 by Lifetree Software Inc. The Mathematical Intelligencer, 8(1):66–71, 1986.

Miller:PROTEXT-IV

[Mil87] John J. H. Miller, editor. PRO-TEXT IV: Proceedings of the Fourth International Conference on Text Processing Systems, 20– 22 October 1987, Boston, USA. Boole Press, Dublin, Ireland, 1987. ISBN 0-906783-80-1 (hardcover), 0-906783-79-8 (paperback). LCCN QA76.9.T48 I571 1987.

Milne:NAMS-37-8-1018

[Mil90] J. S. Milne. Four word processors with TEX capabilities. Notices of the American Mathematical Society, 37(8):1018–1022, October 1990.

Milne:NAMS-39-8-838

[Mil92] J. S. Milne. Three scientific word processors. Notices of the American Mathematical Society, 39(8): 838–840, October 1992. The word processors ChiWriter 4.02, Leo 1.4, and Scientific Word 1.0 are discussed in terms of ease of use, effectiveness in producing TEX code, and other features of interest to users who require TEX output but do not wish to use TEX directly.

Molnar:UMAP-14-2-97-112

[Mol93] Suzanne M. Molnar. A tour of graphical typesetting on the macintosh. The UMAP Journal, 14 (2):97–112, 1993.

Mortenson:RIT91

[Mor91]  Niels G. Mortensen.  Risø-

rapporter i TEX. Matematisk og naturvidenskabelig typografi i praksis (Risø reports in TEX. Mathematical and scientific typography in practice). Risø National Laboratory, Roskilde, Denmark, 1991. ISBN 87-550-1635-9. Prepared using LaTEX, and typeset on a Chelgraph IBX photo-typesetter.

Morin:SUNEXPERT-4-10-42

[Mor93] Richard Morin. Tangled up in TEX. SunExpert Magazine, 4 (10):42–46, October 1993. A brief review of TEX support, including the Comprehensive TEX Archive Network (CTAN), the TEX Users Group, and several books on TEX.

Nadrchal:ESS89

[Nad90]  J. Nadrchal, editor.  Man-

Machine Interface in the Scientific Environment. Proceedings of the 8th European Summer School on Computing Techniques in Physics. Skalský Dvůr, Czecholsovakia, 19–28 September 1989, volume 61 of Computer Physics Communications. European Summer Schools on Computing Techniques in Physics, North-Holland Publishing Company; Elsevier Science Publishers B. V., Amsterdam, The Netherlands, 1990. In addition to the papers presented at the colloquium, each paper is followed by a summary of the discussion about that paper.

Naiman:UGI91

[Nai91]  Avi Naiman.  The Use of

Grayscale for Improved Character Presentation. PhD thesis, University of Toronto, ATTN: Technical Reports, Computer Systems Research Institute, University of Toronto, 6 King’s College Road, Toronto, Ontario M5S 1A1, CANADA, Tel: (416) 978-8751. The cost is CDN$16 if inside Canada, US$16 if outside. Avi’s e-mail address is avi@cvs. rochester.edu., 1991. Technical report CSRI-253.

Nelson:PD78

[Nel78] Roy Paul Nelson. Publication Design. William C. Brown Company Publishers, Dubuque, IA, USA, second edition, 1978. ISBN 0-697-04324-X. 317 (or 295??) pp. LCCN Z246 .N44 1978.

Noot:SPE-13-1-79

[Noo83] Han Noot. Structured text formatting. Software—Practice and Experience, 13(1):79–94, January 1983.

NTG

[NTG ] Nederlandstalige TEX gebruikersgroep, 1989–. Journal of the NTG (Dutch-speaking TEX Users Group).

Oman:CACM-33-5-506

[OC90] Paul W. Oman and Curtis R. Cook. Typographic style is more than cosmetic. Communications of the Association for Computing Machinery, 33(5):506–520, May 1990.

BCS:NAMS-33-1-8

[otB86] Report of the Boston Computer Society. Technical Wordprocessors for the IBM PC and Compatibles. Notices of the American Mathematical Society, 33(1): 8–37, January 1986.

Palais:NAMS-33-1-3

[Pal86] Richard S. Palais. Mathematical text processing. Notices of the American Mathematical Society, 33(1):3–7, January 1986.

Pappas:COMPUTER-23-10-82

[Pap90] T. L. (Frank) Pappas. PostScript output on a LaserJet printer. Computer, 23(10):82–88, October 1990. ISSN 0018-9162.

Pappas:COMPUTER-24-6-92

[Pap91] T. L. (Frank) Pappas. TEX typesetting with a difference. Computer, 24(6):92–93, June 1991. ISSN 0018-9162.

Partl:HTL89

[Par90] Hubert Partl. How to make TEX and LaTEX international. In Nadrchal [Nad90], pages 190– 200. Invited paper.

Pawlowski:LTL93

[Paw93] Wiesław Pawłowski. LaTEX, czyli TEX dla ludzi. broszura z konferencji członków GUSTu, Bachotek, 1993. In Polish.

Popperlier:NTG-92-1-135

[PH92] Nico Poppelier and Amy Hendrickson. Book reviews. Nederlandstalige TEX Geruikersgroup (Dutch-speaking TEX Users Group), 92(1):135–138, May 1992.

Pietrowski:NIPT91

[Pie91] Ken Pietrowski, editor. IS&T’s Seventh International Congress on Advances in Non-impact Printing Technologies, October 6–11, Portland, OR, USA., volume 2. Imaging Science & Technology, ????, 1991.

Plass:CBL82

[PK82] Michael F. Plass and Donald E.

Knuth. Choosing better line breaks, chapter ??, pages 221– 242. Elsevier Science Publishers (North-Holland), Amsterdam, The Netherlands, 1982. ISBN ?? LCCN ??

Popperlier:TB13-2-182

[Pop92] Nico Poppelier. Book reviews:

Jane Hahn, LaTEX for Everyone; Eric van Herwijnen, Practical SGML. TUGboat, 13(2):182– 184, July 1992.

Wood:plj

[POS] PostScript Language Journal.

Pipeline Associates, Inc., P. O. Box 5763, Parsippany, NJ 07054.

Partl:DL93

[PSH93] Hubert Partl, Elis. Schlegl, and Irene Hyna. Una descripción de LaTEX, 1993. Spanish translation by Tomas Bautista of German introduction to LaTEX, LaTEX-Kurzbeschreibung. Includes material on the spanish style option. Available in Spain from

ftp.eunet.es (also known as

goya.uu.es) in the directory

info/unix/text/TeX/spanish/

tutorial, as well as from the main TEX archives.

Rahtz:CSTR-89-7

[Rah89] Sebastian Rahtz. A survey of TEX and graphics. Technical Report CSTR 89-7, University of Southampton, Department of Electronics and Computer Science, Southampton SO9 5NH, UK, October 1989.

Adobe:PLP88

[Rei88] Glenn C. Reid. PostScript Language Program Design. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-14396-8. xii + 224 pp. LCCN QA76.73.P67 R45 1988.

Rice:BDT78

[Ric78] Stanley Rice. Book Design—Text Format Models. R. R. Bowker Company, 1180 Avenue of the Americas, New York, NY 10036, USA, 1978. ISBN 0-8352-1045-6. xiv + 215 pp. LCCN Z245 R5 1978. US$17.50.

Rosson:CACM-31-11-1288

[RKM88] Mary B. Rosson, Wendy Kellogg, and Susanne Maass. The designer as user: building requirements for design tools from design practice. Communications of the Association for Computing Machinery, 31(11):1288–1298, November 1988. From Computing Reviews: “InterViews objects are combined through a model similar to the boxes and glues in Knuth’s TEX, which allows interface objects to be combined with little concern for undesired interactions.”.

Rolland:LGP93

[Rol93] Christian Rolland. LaTEX guide pratique. Addison-Wesley, Reading, MA, USA, July 1993. ISBN 2-87908-025-8. 280 pp. FF 190. Distributed by Bourdas, code 009808.

Rosset:CIRCE-65

[Ros84] J.-L. Rosset, editor. S.A.PR.I.S.T.I.

(système assisté de PRoduction intégrant simplement textes et images). Supplément au No. 65, CIRCE, Orsay, France, 1984.

Roth:RWP88

[Rot88] Stephen E. Roth, editor. Real World PostScript. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-06663-7.

Roy:TW84

[Roy84] Yves Roy. TEX/WEB, et le traitement de textes mathématiques. Masson Editeur, Paris, France, 1984.

Romberger:TEX85-27

[RS85] Staffan Romberger and Yngve Sundblad. Adapting TEX to Languages that use Latin Alphabetic Characters. In Lucarella [Luc85b], pages 27–40. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

Rubinstein:DTI88

[Rub88]  Richard Rubinstein.  Digital

Typography—An Introduction to Type and Composition for Computer System Design. Addison-Wesley, Reading, MA, USA, November 1988. ISBN 0-201-17633-5. xi + 340 pp. LCCN Z253.3 .R8 1988. US$32.25. Reprinted with corrections.

Ruder:TG88

[Rud88] Emil Ruder. Typographie; Ein Gestaltungsbuch. Niggli/Hatje, Heiden, Stuttgart, Germany, fifth revised edition, 1988. Contains complete English and French translation.

Ruggles:STAN-CS-83-971

[Rug83] Lynn Ruggles. Letterform design systems. Technical Report STAN-CS-83-971, Stanford University, April 1983.

Rycko:CNA93

[Ryć93]  Marek Ryćko.  Co nieco o

akapicie. broszura z konferencji członków GUSTu, Bachotek, 1993. In Polish.

Samuel:STAN-CS-83-985

[Sam83] Arthur Samuel. First grade TEX. A beginner’s TEX manual. Technical Report STAN-CS-83-985, Stanford University, November 1983. Also published by TUG as TEXniques No. 11.

Sutton:TB90

[SB90] James Sutton and Alan Bartram.

Typefaces for Books. The British Library and New Amsterdam Books, London, UK and New York, NY, USA, 1990. ISBN 1-56131-016-6, 1-56131-017-4 (paperback). 288 pp. LCCN Z250 .S896 1990. Contains full-page type samples in a variety of point sizes for a large number of typefaces.

Schwartz:WYA88

[Sch88a] Laurens R. Schwartz. What You Aren’t Supposed To Know About Writing And Publishing: An Expos/’e of Editors, Agents, Publishing Houses, and More — An Insider’s Report. Shapolsky Publishers, New York, NY, USA, 1988. ISBN 0-944007-02-3. viii + 173 pp. LCCN KF300.Z9 S35 1988.

Schwarz:ET88

[Sch88b] Norbert Schwarz. Einführung in TEX. Addison-Wesley Verlag, Bonn, Germany, second edition, 1988. ISBN 3-925118-97-7. Available in English translation, [KS90], and in Dutch translation, [Sch90].

Schumann:PBT89

[Sch89]  Lothar Schumann.  Profes-

sioneller Buchsatz mit TEX. R. Oldenbourg Verlag, Munich, Germany and Vienna, Austria, 1989. ISBN 3-486-21173-0.

Schwarz:IT90

[Sch90] Norbert Schwarz. Inleiding TEX. Addison-Wesley Europe, Amsterdam, The Netherlands, 1990. ISBN 90-6789-151-7. Dutch translation of [Sch88b]. See also the English translation, [KS90].

Schrod:CG-10-81

[Sch91a]  Joachim Schrod.  An inter-

national version of makeindex. Cahiers GUTenberg, (10–11):81– 90, September 1991.

Schumann:PBT91

[Sch91b] Lothar Schumann. Professioneller Buchsatz mit TEX: Lehrbuch für Anwender. R. Oldenbourg Verlag, Munich, Germany and Vienna, Austria, 1991. ISBN 3-486-21736-4.

Schrell:DTS93

[Sch93] DANTE-TEX-stammtisch in wuppertal: Jahrbuch 1992/93, 1993.

Scherber:OST94

[Sch94] Peter Scherber. Offizin—Schriftenreihe zu TEX, LaTEX, und Metafont (Band 1, Hrsg. von DANTE e.V). Addison-Wesley, Reading, MA, USA, 1994. ISBN 3-89319-396-0. 190 pp. DM 49.90.

Seroul:PLT89

[Ser89] Raymond Seroul. Le petit Livre de TEX. InterEditions, Paris, France, 1989. ISBN 2-7296-0233-X. 317 pp.

Sewell:WPL89

[Sew89] E. Wayne Sewell. Weaving a Program: Literate Programming in WEB. Van Nostrand Reinhold, New York, NY, USA, 1989. ISBN 0-442-31946-0. xx + 556 pp. LCCN QA76.73.W24 S491 1989. US$34.95.

Skillin:WIT74

[SG74] Marjorie E. Skillin and Robert Malcolm Gay. Words into Type. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 1974. ISBN 0-13-964262-5. xx + 583 pp. LCCN PN160.S52 1974. US$39.95.

Shultis:LN94

[Shu94]  J. Kenneth Shultis.  LaTEX

Notes: Practical Tips for Preparing Technical Documents. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-120973-6. xi + 180 pp. LCCN Z253.4.L38 S48 1994. US$32.00.

Siegel:euler-proj

[Sie85] David R. Siegel. The Euler project at Stanford. Technical report, Stanford University, 1985.

Siebenmann:NAMS-33-4-597

[Sie86] L. C. Siebenmann. Toward Wider Use of TEX Typesetting. A Pre–TEX Manifesto. Notices of the American Mathematical Society, 33(4):597–607, August 1986.

Siemoneit:TG89

[Sie89] Manfred Siemoneit. Typographisches Gestalten. Polygraph Verlag, Frankfurt am Main, Germany, second edition, 1989.

Sawyer:TPS93

[SK93] Stanley A. Sawyer and Steven Krantz. A TEX Primer for Scientists. CRC Publishers, 2000 Corporate Blvd., Boca Raton, FL 33431, USA, 1993. ISBN 0-8493-7159-7. 175 pp. US$29.95. To appear March 1993.

Seroul:BBT91

[SL91] Raymond Seroul and Silvio Levy. A Beginner’s Book of TEX. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97562-4, 3-540-97562-4. xii + 282 pp. LCCN Z253.4.T47 S47 1991. This is a translation and adaption by Silvio Levy of [Ser89].

Snow:TB92

[Sno92] Wynter Snow. TEX for the Beginner. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-54799-6. xii + 377 pp. LCCN Z253.4.T47 S64 1992.

Southall:INRIA-133-85

[Sou85a] Richard Southall.

and the problem of type design. In André [And85], pages 133– 145.

Southall:STAN-CS-85-1074

[Sou85b] Richard Southall. Designing new typefaces with Metafont. Technical Report STAN-CS-85-1074, Stanford University, September 1985.

Schwartz:LSD92

[SP92] Stefan Schwartz and Rudolf Potuček. LaTEX—Satzkunst statt DTP (LaTEX—The Art of Typography rather than DTP). Vogel Verlag, Postfach 67 40, D-8700 Würzburg, Germany, 1992. ISBN 3-8023-1178-7. DM34.00.

Spivak:PTM85

[Spi85] Michael D. Spivak. PCTEX Manual. 12 Madrona Street, Mill Valley, CA 94941, USA, 1985.

Spivak:JTG86

[Spi86a] Michael D. Spivak. The Joy of TEX—A Gourmet Guide to Typesetting with the AMS-TEX macro package. American Mathematical Society, Providence, RI, USA, 1986. ISBN 0-8218-2999-8. xviii + 290 pp. LCCN Z253.4.T47 S6731 1986.

Spivak:MPT86

[Spi86b] Michael D. Spivak. Manual “PCTEX” ou TEX en 9 leçons. Orsay, France, 1986. French translation of [Spi85].

Spivak:LS89

[Spi89] Michael D. Spivak. LAMS-TEX, The Synthesis. The TEXplorators Corporation, 3701 W. Alabama, Suite 450-273, Houston, TX 77027, USA, 1989. viii + 289 pp.

Spivak:JTG90

[Spi90] Michael D. Spivak. The Joy of TEX—A Gourmet Guide to Typesetting with the AMS-TEX macro package. American Mathematical Society, Providence, RI, USA, 2nd revised edition, 1990. ISBN 0-8218-2997-1. xxii + 309 pp. LCCN Z253.4.T47 S6731 1990. US$40.00.

Spivak:LTW91

[Spi91] Michael D. Spivak. The L^{A}MS-TEX Wizard’s Manual. The TEXplorators Corporation, 3701 W. Alabama, Suite 450-273, Houston, TX 77027, USA, 1991.

Smith:SIGPLAN-26-1-14

[SS91] Lisa M. C. Smith and Mansur H. Samadzadeh. An annotated bibliography of literate programming. ACM SIGPLAN Notices, 26:14–20, January 1991.

Tang:STAN-CS-81-848

[sT81] Chih sung Tang. On the problem of inputting Chinese characters. Report No. STAN-CS-81-848, Stanford University, Department of Computer Science, 1981.

Steen:UMAP-2-3-17-35

[Ste81] Lynn Arthur Steen. Technical text editing. The UMAP Journal, 2(3):17–35, 1981.

Stone:COMPUTER-25-12-46

[Sto92] Harold S. Stone. Copyrights and author responsibilities. Computer, 25(12):46–51, December 1992. This paper discusses U. S. copyright law and its implications for the reuse of published text. Guidelines are given for the citation and quoting of such material.

Steemers:RRG-26

[SvdL89] L. Steemers and C. G. van der Laan. Journal style guidelines. RC-Rapport 26, Rekencentrum der Rijksuniversiteit te Groningen, Groningen, The Netherlands, 1989.

Swanson:MTC86

[Swa86] Ellen Swanson. Mathematics into Type; Copy Editing and Proofreading of Mathematics for Editorial Assistants and Authors. American Mathematical Society, Providence, RI, USA, revised edition, 1986. ISBN 0-8218-0053-1. x + 90 pp. LCCN T11.4.S96 1979.

Tatarkiewicz:Komputer-8-36-37

[Tat88] Jakub Tatarkiewicz. TEXnika ładnego druku. Computer, (8): 36–37, 1988. In Polish.

TEXEURO

[TEX a] TEX-euro, 1989–. This is an unmoderated discussion list for TEX with emphasis on European issues. To subscribe, send a request with the text subscribe

tex-euro to listserv@dhdurz1.
bitnet.

TEXHAX

[TEX b] TEXhax, 1987–. This is a TUGsupported moderated electronic mailing list. To subscribe, send

a request to texhax-request@
june.cs.washington.edu.

Clark:texline

[TEX c] Malcolm Clark, editor. TEXline, 1987–. ISSN 0961-3978. This is an informal newsletter of the TEX community.

TEXMAG

[TEX d] TEXmag, 1988–. This is an electronic magazine with articles about TEX. To subscribe, send a request with the text subscribe

                                        texmag-l to listserv@vm.byu.
                                        edu.

Tobin:elements-mf-style

[Tob85] Georgia K. M. Tobin. The elements of metafont style, 1985.

Trickey:SIGACT-19-2-55

[Tri88] Howard Trickey. Using LaTEX. ACM SIGACT News, 19(2):55– 57, Summer 1988.

Tschichold:TM86

[Tsc86] Jan Tschichold. Typographische Mitteilungen, Sonderheft Elementare Typographie. Universitätsdruckerei und Verlag Hermann Schmidt, Mainz, Germany, 1986. ISBN 3-87439-233-3. 36 pp. DM24.

Tschichold:AAF87

[Tsc87] Jan Tschichold. Ausgewählte Aufsätze über Fragen der Gestalt des Buches. Birkhäuser, Basel, Switzerland, 1987. Second printing.

Tschichold:NT91

[Tsc91a] Jan Tschichold. Die Neue Typographie. Universitätsdruckerei und Verlag Hermann Schmidt, Mainz, Germany, 1991. 240 pp. DM60.

Tschichold:FB91

[Tsc91b] Jan Tschichold. The Form of the Book. Lund Humphries, London, 1991. ISBN 0-85331-623-6. xviii + 180 pp. LCCN Z116.A3 T7613 1991b. UK£18.95. Originally published in German as Ausgewählte Ausfsätze über Fragen der Gestalt des Buches und der Typographie by Birkhäuser Verlag, Basel, 1975.

Tufte:VDQ83

[Tuf83] Edward R. Tufte. The Visual Display of Quantitative Information. Graphics Press, Box 430, Cheshire, CT 06410, USA, 1983. ISBN 0-9613921-0-X. 197 pp. LCCN QA90 .T83 1983. US$34.00.

Tufte:EI90

[Tuf90] Edward R. Tufte. Envisioning Information. Graphics Press, Box 430, Cheshire, CT 06410, USA, 1990. 126 pp. LCCN QA90 .T82 1990. US$48.00.

Tung:STAN-CS-80-824

[Tun80] Tung Yun Mei. LCCD, A Language for Chinese Character Design. Report No. STAN-CS-80-824, Stanford University, Department of Computer Science, 1980. See also Software—Practice and ExperienceVol. 11 (1981), No. 12, pp. 1273–1292.

Tymms:AI87

[TW87] W. R. Tymms and M. D. Wyatt. The Art of Illuminating. Chartwell Books, Inc., 110 Enterprise Avenue, Secaucus, New Jersey 07094, USA, 1987. ISBN 1-55521-198-4. 104 pp. LCCN ND3310 .T95 1987.

UKTEX

[UKT ] UKTEX, 1987–. This is an electronic discussion list for TEX issues in the United Kingdom.

Ulichney:DH87

[Uli87] Robert Ulichney. Digital Halftoning. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-21009-6. xiv + 362 pp. LCCN T385 .U451 1987.

Urban:IL86

[Urb86] Michael Urban. An Introduction to LaTEX. Prepared for the TRW Software Productivity Project; reprinted with permission and distributed by TUG, 1986.

Urban:PPL91

[Urb91] Michael Urban (adaptation française de Éric Cornelis). Premiers pas en LaTEX. Cahiers GUTenberg, May 1991. French translation of [Urb86].

Utting:TEX85-183

[Utt85] Ian Utting. The Rôle of Device Independent Output. In Lucarella [Luc85b], pages 183– 191. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

VanWyk:STAN-CS-80-803

[Van80] C. J. Van Wyk. A Language for Typesetting Graphics. Report No. STAN-CS-80-803, Stanford University, Department of Computer Science, 1980.

vonBechtolsheim:TP92

[vB92a] Stephan von Bechtolsheim. TEX in Practice. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97296-X. 1587, with 31 illustrations pp. LCCN Z253.4.T47 B4 1993. US$169.00. Four volume set.

vonBechtolsheim:TPB92

[vB92b] Stephan von Bechtolsheim. TEX in Practice: Basics, volume 1. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97595-0. 359, with 9 illustrations pp. LCCN Z253.4.T47 B4 1993. US$49.00.

vonBechtolsheim:TPP92

[vB92c] Stephan von Bechtolsheim. TEX in Practice: Paragraphs, Math, and Fonts, volume 2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97596-9. 384 pp. LCCN Z253.4.T47 B4 1993. US$49.00.

vonBechtolsheim:TPT92

[vB92d] Stephan von Bechtolsheim. TEX in Practice: Tokens, Macros, volume 3. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97597-7. 544, with 22 illustrations pp. LCCN Z253.4.T47 B4 1993. US$49.00.

vonBechtolsheim:TPO92

[vB92e] Stephan von Bechtolsheim. TEX in Practice: Output Routines, volume 4. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97598-5. 300 pp. LCCN Z253.4.T47 B4 1993. US$49.00.

Bechtolsheim:TP93a

[vB93a] Stephan von Bechtolsheim. TEX in Practice: Basics, volume 1. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-97595-0, 3-540-97595-0. LCCN Z253.4.T47 B4 1993.

Bechtolsheim:TP93d

[vB93b] Stephan von Bechtolsheim. TEX in Practice: Output Routines, Tables, volume 4. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-97598-5, 3-540-97598-5. LCCN Z253.4.T47 B4 1993.

Bechtolsheim:TP93b

[vB93c] Stephan von Bechtolsheim. TEX in Practice: Paragraphs, Math, and Fonts, volume 2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-97596-9, 3-540-97596-9. LCCN Z253.4.T47 B4 1993.

Bechtolsheim:TP93c

[vB93d] Stephan von Bechtolsheim. TEX in Practice: Tokens, Macros, volume 3. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-97597-7, 3-540-97597-7. LCCN Z253.4.T47 B4 1993.

vanderLaan:RRG-24

[vdLCL89] C. G. van der Laan, D. C. Coleman, and J. R. Luyten. SGML– LaTEX 1. Mathematical formulas. RC-Rapport 24, Rekencentrum der Rijksuniversiteit te Groningen, Groningen, The Netherlands, 1989. (English version).

vanderLaan:RRG-27

[vdLL89] C. G. van der Laan and J. R.

Luyten. LaTEX templates: Letter and article. RC-Rapport 27, Rekencentrum der Rijksuniversiteit te Groningen, Groningen, The Netherlands, November 1989.

vanderVeer:HCI88

[vdVM88] Gerrit C. van der Veer and Gijsbertus Mulder, editors. Humancomputer Interaction: Psychonomic Aspects. Conference of the Dutch Psychonomic Society. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-18901-7.

vanHuu:TEX85-67

[vH85] Le van Huu. TEX and ISO/STPL standard. In Lucarella [Luc85b], pages 67–78. ISBN 0-201-13399-7. LCCN Q224.E97 1985.

vanHerwijnen:PS90

[vH90]  Eric van Herwijnen.  Prac-

tical SGML. Kluwer Academic Publishers Group, Norwell, MA, USA, 1990. ISBN 0-7923-0635-X. xviii + 307 pp.

LCCN QA76.73.S44 V36 1990. UK£24.90, US$49.00.

Vulis:MTA92

[Vul92] Michael Vulis. Modern TEX and its Applications. CRC Publishers, 2000 Corporate Blvd., Boca Raton, FL 33431, USA, 1992. ISBN 0-8493-4431-X. 275 pp. LCCN Z253.4.T47V84 1993. US$32.95.

vanVliet:TPDM86

[vV86] J. C. van Vliet, editor. Text Processing and Document Manipulation: Proceedings of the International Conference, University of Nottingham, 14–16 April 1986. Cambridge University Press, New York, April 1986. ISBN 0-521-32592-7. LCCN QA76.9.T48 T481 1986.

vanVliet:DMT88

[vV88] J. C. van Vliet, editor. Document Manipulation and Typography. Proceedings of the International Conference on Electronic Publishing, Document Manipulation and Typography, Nice (France), April 20–22, 1988. Cambridge University Press, New York, 1988. ISBN 0-521-36294-6. LCCN Z286.E43 I571 1988. US$47.50.

Walden:MFF87

[Wal87a] Jeff Walden. More File Formats for Popular PC software: A Programmer’s Reference. Wiley, Chichester, UK, 1987. ISBN 0-471-85077-2. 369 pp. LCCN QA76.9.F5 W35 1987. US$24.95.

From Computing Reviews: “The volume under review deals with (among others) WordPerfect, SuperCalc 4, and the unusual Microsoft Rich Text Format (RTF). RTF has the distinction of being entirely readable ASCII and is reminiscent of TEX command format.” Conversion software from RTF to LaTEX written by Daniel Taupin is available from major TEX archive sites.

Waldschmidt:CGF-6-4-359

[Wal87b] H. Waldschmidt. An editor for constructing graphics with TEX. Computer Graphics Forum, 6(4): 359–364, December 1987.

Walsh:MTW94

[Wal94] Norman Walsh. Making TEX Work. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1994. ISBN 1-56592-051-1. xxxvi + 483 pp. US$29.95.

Wolcott:NBS-424

[WH76] N. M. Wolcott and J. Hilsenrath. A contribution to computer typesetting techniques: Tables of coordinates for Hershey’s repertory of occidental type fonts and graphic symbols. Technical Report Special Publication 424, U. S. National Bureau of Standards, April 1976.

White:MG83

[Whi83] Jan V. White. Mastering Graphics. R. R. Bowker Company, 1180 Avenue of the Americas, New York, NY 10036, USA, 1983.

ISBN 0-8352-1704-3. xiii + 180 pp. LCCN Z253 .W47 1983.

White:HST87

[Whi87]  Alex White.  How to spec

type. Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 1987. ISBN 0-8230-2473-3. 128 pp. LCCN Z253.W46 1987. US$14.95.

White:GDE88

[Whi88] Jan V. White. Graphic Design for the Electronic Age. Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 1988. ISBN 0-8230-2122-X. xi + 211 pp. LCCN Z286.D47 W5 1988.

White:GP90

[Whi90] Jan V. White. Great Pages. Serif Publishing, A subsidiary of Xerox Corporation, 701 South Aviation Boulevard, El Segundo, CA 90245, USA, 1990. ISBN 1-878567-01-2. 128 pp. LCCN Z286.D47 W5 1990.

Wieckowski:PCkurier-3-15-89

[Wiȩ92] Z. Wiȩckowski. TEX — polskie znaki inaczej. PCkurier, (15):89, lipiec 1992. In Polish.

Williamson:MBD83

[Wil83] Hugh Williamson. Methods of Book Design: The Practice of an Industrial Craft. Yale University Press, New Haven, CT, USA, third edition, 1983. ISBN 0-300-02663-3 (hardcover), 0-300-03035-5 (paperback). xiv + 391 pp. LCCN Z116.A3 W5 1983.

UK£13. Bibliography, many examples and illustrations, index.

Winkel:COLL-MICRO-5-1-93

[Win88] Brian J. Winkel. TEX: The software used to produce this journal. Collegiate Microcomputer, 5 (1):93–96, April/May 1988.

Abrahams:TI90

[wKBH90] Paul W. Abrahams with Karl Berry and Kathryn A. Hargreaves. TEX for the Impatient. Addison-Wes-ley, Reading, MA, USA, 1990. ISBN 0-201-51375-7. xvii + 357 pp. LCCN Z253.4.T47 A27 1990.

Abrahams:TI91

[wKBH91] Paul W. Abrahams with Karl Berry and Kathryn A. Hargreaves. TEX per l’impaziente. Addison-Wes-ley, Reading, MA, USA, 1991. ISBN 88-7192-022-8. 396 pp. This is an translation to Italian of [wKBH90], with additional material on Italian hyphenation.

Weyerich:LLL93a

[WKS^{+}93a] Astrid Weyerich, Rainer Klötzer, Sabine Stromp, Kerstin Rieder, and Rainer Oesterreich. LaTEX Lernen leicht gemacht—Eine aufgabenorientierte Anleitung— Teil A: Arbeitsheft. Technische Universität Berlin, Institut für Humanwissenschaft in Arbeit und Ausbildung, 10623 Berlin, Germany, 1993. ISBN 3-7983-1546-9. 100 pp.

Weyerich:LLL93b

[WKS^{+}93b] Astrid Weyerich, Rainer Klötzer, Sabine Stromp, Kerstin Rieder, and Rainer Oesterreich. LaTEX Lernen leicht gemacht—Eine aufgabenorientierte Anleitung— Teil B: Manual. Technische Universität Berlin, Institut für Humanwissenschaft in Arbeit und Ausbildung, 10623 Berlin, Germany, 1993. ISBN 3-7983-1546-9. 176 pp.

Wonneberger:L87

[Won87] Reinhard Wonneberger. LaTEX. Addison-Wesley Kompaktführer. Addison-Wesley Verlag, Bonn, Germany, 1987. ISBN 3-925118-46-2. 141 pp.

Wonneberger:L88

[Won88] Reinhard Wonneberger. LaTEX. Addison-Wesley Kompaktführer. Addison-Wesley Verlag, Bonn, Germany, second edition, 1988. ISBN 3-89319-152-6. xiv + 141 pp.

Wonneberger:SDP89

[Won90] Reinhard Wonneberger. Structured document processing: the LaTEX approach. In Nadrchal [Nad90], pages 177–189. Invited paper.

Wonneberger:TPT92

[Won92] Reinhard Wonneberger. TEX for philological typesetting. In Proceedings of the XXXII International Congress for Asian and North African Studies. Hamburg, 25th–30th August 1986, number IX in Zeitschrift der Deutschen Morgenländischen Gesellschaft, Supplement, pages 595–596. Steiner, Stuttgart, Germany, 1992. ISBN 3-515-04808-1.

Wonneberger:L93

[Won93] Reinhard Wonneberger. LaTEX.

Addison-Wesley Kompaktführer. Addison-Wesley Verlag, Bonn, Germany, third edition, April 1993. ISBN 3-89319-589-0. xvi + 166 pp. DM29.80.

Wyrostek:LSP92

[Wyr92] Piotr Wyrostek, translator.

LaTEX: System przygotowywania dokumentów. Przewodnik użytkownika i podrȩcznik. Ariel Sp. z o.o, Kraków, Poland, 1992. ISBN 83-900460-1-6. 91000 zl. Polish translation of [Lam85].

Zapf:HZD87

[Zap87] Hermann Zapf. Hermann Zapf and His Design Philosophy: Selected Articles and Lectures on Calligraphy and Contemporary Developments in Type Design, with Illustrations and Bibliographical Notes, and a Complete List of his Typefaces. Society of Typographic Arts, Chicago, IL, USA, 1987. ISBN 0-941447-00-6. 254 pp. LCCN Z250 .Z323 1987. US$45.00.

Zlatuska:ET92

[Zla92]  Jiří Zlatuška, editor.  Eu-

roTEX 92: Proceedings of the 7th European TEX Conference. Masarykova Universita, Brno, Czechoslovakia, September 1992. ISBN 80-210-0480-0.